A Complete Bibliography of Proceedings of the ACM on Measurement and Analysis of Computing Systems

(POMACS)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

20 August 2024 Version 1.04

Title word cross-reference

1/N [Gas17]. 3 [ADGF^{+}24, KKRK22, LKJK20, LZGN22, LGC^{+}18]. $9.81 [FS20]. D [YXL21, HV21]. k [CKMP19, SGHB20, WHBW22].

-Accurate [Gas17]. -Core [CKMP19]. -hops [YXL21].

1 [CP18, SHBSW20, YS24].

2021 [CCR21]. 2022 [AGU22, CCR22a, CCR22b]. 2023 [AGU23a, AGU23b, CFR23]. 2024 [CFR24a, CFR24b]. 256 [BRC21].

5G [PLX22]. 5G-LTE [PLX22].

6 [LC23]. 60 [ANBK22].

Accelerated [KKS22]. Accelerating [Cas17, LJH^{+}23]. Accelerator [KKP^{+}22]. Access [LYE19, VPK18, WM20, WXW22]. Access-Delay [LYE19]. Accesses [DMB^{+}22]. Accessibility [SCH^{+}24]. Accurate [BNS17, Gas17, JHWC21, LPF^{+}22]. Achievable [RB20]. Achieving [WW20, YCX^{+}19]. Acoustic [IYRR18]. Acquisition [MLW^{+}23]. Across [SATN22].

Action [LC23]. Activity [LTS23]. Ad [SVL20, TMM20]. Adaptation [WCX^{+}18]. Adaptive [LPJ23a, SBY19, ZCM^{+}22]. Additional [YPA^{+}18]. additive [WS17a]. Addresses [MLHC20]. Adoption [AZG24]. adPerf [PBLC21]. Ads [PBLC21]. Adversarial [CSX17, HLH23, LSNI21, LM18, YHW23]. Adversary [RKTV22]. Adversary-resistant [RKTV22]. Advertising [PPL^{+}21, YWB19]. Age [SHBSW18]. Age-Based [SHBSW18]. Agent [LVW19, SGS19, VSS22, ZQX^{+}23, ALZ22]. Agents [BG23]. Aggregate [PTD20]. Aggregation [KLCS19, LCS^{+}21]. Aggressive [LZC^{+}23]. Agnostic [AASS18]. Agreements [WM20]. Ahead [SLJ19]. AI [ILL^{+}23]. Algorithm [DCKG19, JIAM19, TTL^{+}21, YWH17, YDH^{+}18]. Algorithmic [NT19, VS19]. Algorithms [CMM22, CYC^{+}19, CHT^{+}22, CCG21, FKP^{+}21, HCZ^{+}21, KMM21, LNY21, LCZ^{+}23, MSPI22, SZL^{+}20, YZH^{+}21, YCH^{+}22, ZWT^{+}17]. Alignment [CKMP19, DCKG19]. All-Flash [ARK^{+}22]. Allocation [AIAR^{+}24, ILL^{+}23, RKTV22, SIN22, TSLG^{+}20, YLIR24, ZLW17]. Allocations [TP18]. Alternative [FS19]. Ambivalence [ZMC24]. Among [PJDQ20]. amplify [ZMC24]. Analysis [AASS18, ALR17, BRC21, BNS17, CYG^{+}17, CMH^{+}17, CF23, CLE^{+}24, DCKG19, DSR24, HBV19, HSV19, LKS^{+}17, LML^{+}20, MMS^{+}23, NVGT20, OKB^{+}18, PSC19, RSJ24, SAM^{+}22, SHBSW18, SSM^{+}17, WM20, YHG^{+}17]. Analytics [AEKB20, KLCS19]. Analyzing [YWB19]. Android [WWH^{+}22, ZYH^{+}19]. Anonymity [FVB^{+}18, GFS23, VFV17]. App [ZYH^{+}19]. Application [DDT^{+}22, LCZ^{+}23, SZL^{+}20, YNJS21, ZYH^{+}19, LJH^{+}23]. Applications [BZY^{+}22, LZGN22]. Apportionment [HLP23]. Approach [AdV18, AGS17, BCM17, FZCL18, GTL^{+}17, KLCS19, LMSC22, NT19, SSM^{+}17, TSLG^{+}20, WKWY19, ZZLL21]. Approximate [KKT^{+}19, PHSB21, WLS19]. Approximation [DE20, Gas17, GV17, HCZ^{+}21, Zho22]. Approximations [AG22, GV24, RBT21, Yin17]. Apps [LZC^{+}23, ZTY^{+}21]. Arbitrage [LGL23]. Architectural [GHT^{+}23]. Architecture [ARK^{+}22, KKT^{+}19, LPF^{+}22, LKJK20, LYJ^{+}21]. Architecture-Aware [KKT^{+}19]. Architectures [GFL^{+}22]. Argus [RS22]. Armed [SGS19, YPA^{+}18]. Arms [CPF20]. Assessment [NT19]. Assisted [GKK20]. Asymptotic [Cas17, CMM22, DV18, KV21, SCS22, Van22, WW20]. Asymptotically [MBvL18]. Asynchronous [LPJ23b]. At-Scale [CLE^{+}24]. Attacker [WS17a]. Attacks [FM19, HLH23, IYRR18]. Attraction [Van19]. Auctions [KMM21, TMM20]. Augmentation [JLGJ^{+}24, RS22]. Automata [LPJ23b]. Automated [ILL^{+}23]. Automatic [SMF^{+}22]. Autonomous [ZHZ^{+}21]. Autonomy [BG23]. Aware [AMBK23, BSA^{+}21, KKT^{+}19, KKRK22, LCZ^{+}23, LYJ^{+}21, VPK18, XGS24]. Away [CWW^{+}23]. Axiomatizing [ZMSS19].

Back [AZG24]. Backbone [SAM^{+}22]. Backend [ILL^{+}23]. Backhaul [YHG^{+}17]. Balanced [BCM17, LGH^{+}17, PJDQ20, YLX17]. Balancing [ALR17, GSHB19, HBV19, HV21, HSV19, MCS^{+}19, MBvL18, TS19, WZS20, ZWT^{+}17, ZWT^{+}18, ZTS18, ZLX^{+}23]. Bandit [CES19, YPA^{+}18, ZZLL21]. Bandits [CPF20, CCG21, DV18, LZJ23, SGS19, VSS22]. Bandwidth [XVK22, ADGF^{+}24]. Bankrupting [PPL^{+}21]. Barrier [SCY23]. Based [AdV18, NGB^{+}21, SHBSW18, TS19, YWB19, ZLL^{+}23, ZZCC17, BKO^{+}23, CMH^{+}17, CAJK17, FM19, KLCS19, SAM^{+}22, Van19, ZTS18, PYL24]. Batch [AI19]. Batching [GFS23]. Beamforming [ZSPZ23]. Behavior [CMM22, HWW^{+}20]. Behind [ASV^{+}21]. Below [HSV19]. Benchmark [HHP20, WZT19]. Best [SZL19]. Better [CJC^{+}19, GTL^{+}17, GAES23]. between [WM20]. Beyond [YLIR24]. BGP [AZG24, SMF^{+}22]. Bias [AG23]. Biased [LZJ23]. Bidding [KMM21, TMM20]. Big [ZZCC17]. Billion [KCT18, NNVD17]. Billion-Scale [KCT18, NNVD17]. Bin [HXW23, LT22]. Bin-Packing [HXW23]. Bitcoin [VFV17]. Bloated [ZAA^{+}24]. Block [ZLX^{+}23]. Blockchain [GSWV20, HWW^{+}20, TKFJ23]. Blockchains [DAR^{+}21, RKTV22]. Blocking [SVL20]. Bloom [DSR24, MBLYR23]. BONES [WSC^{+}24]. Bonus [SCYO21]. Bootstrapped [EC18]. Bot [LGL23]. Bottleneck [RGBY^{+}19, ZMW^{+}20]. Bottlenecks [RT24]. Bought [FS20]. Bound [CP18]. Bounded [KCT18, SZL19]. Bounds [BBHB18]. Bregman [AN20]. Bregman-style [AN20]. Browser [WLZ^{+}22]. Browsers [PSC19]. Budgets [YLIR24]. Buffer [BKO^{+}23, CPR19, YWH17]. Buffers [AAS23]. Burst [WMMC23]. Burstable [WUNK17]. Byzantine [CSX17].

Cables [RJ23]. Cache [QTES19, WXW22, ZKAV20, AMA18]. Caching [AMA18, AGS17, BBHB18, BBS20, GAES23, LSNI21, LYE19, MSPI22, TQJS18]. Cake [KB20]. Calculate [JIAM19]. Calculus [BNS17, BNS22]. Calibrated [LYR22]. Canonical [DCKG19]. Capabilities [WZT19]. capacity [ADGF^{+}24]. capital [ZMC24]. Capsules [NB23]. Capturing [NB23]. Carbon [HLB^{+}23, LCZ^{+}23]. Carbon-Aware [LCZ^{+}23]. Carbon-Efficiency [HLB^{+}23]. CarbonScaler [HLB^{+}23]. Carlo [LKE19]. Cartography [RJ23]. Cascade [NNVD17]. Cascades [HC19]. Case [FWB17, GKK20, KKRK22, XVK22, YLIR24]. Catchment [SK19]. Causal [PSC19, ZL18]. Cause [LML^{+}20]. Cellular [XJ22, YHG^{+}17]. Censored [AS17]. Censorship [JHWC21]. Census [MSJ^{+}19]. Centaur [LKJK20]. Center [PZB^{+}23]. Centers [AAJV20, IYRR18, ZCW21]. Centralization [KALB23]. Centric [PJDQ20]. Cerberus [GZB^{+}21]. Chain [CT19]. Change [LPJ23a]. Channel [AMBK23, IYRR18, PSS23, TWFO20]. Characterization [BZY^{+}22, CYG^{+}17, CMM22, CLE^{+}24, LKS^{+}17, MYS^{+}22]. Characterizing [CLS^{+}21, KKS22, LGL23, PBLC21, SvKB^{+}20, WLdA^{+}17, WLZ^{+}22, XWG^{+}21, ZMW^{+}20]. Charging [SZL^{+}20]. Chasm [ZHM^{+}21]. China [ZMW^{+}20]. Chipotle [FS20]. Chips [LKS^{+}17]. Choices [GZB^{+}21, HV18]. Circuit [ZCW21]. Class [CAJK17, JNTG18]. Classification [ASV^{+}21, CNK^{+}23]. Clean [SHBSW18]. Clocks [TL20]. Close [TLL^{+}18]. Closed [Cas17]. Closer [SCH^{+}24]. Cloud [CYC^{+}19, HLB^{+}23, IKS21, PJDQ20, PG17, SIS17, WUNK17, ZLW17, ZLX^{+}23]. Cloud-Centric [PJDQ20]. Clouds [AZW^{+}23, BKO^{+}23, DDT^{+}22]. CoBF [ZSPZ23]. Codes [MCS^{+}19]. Coding [WLS19]. Collaboratively [SZL19]. Collapse [CBvL24]. Collection [ASME19]. Combinatorial [CCG21]. Come [Van22]. Communication [DBS17, FKP^{+}21, LVW19]. Communities [SMF^{+}22]. Compact [SQ20]. Comparative [MMS^{+}23, PLX22]. Competitive [LYP^{+}19, LMSC22, SLJ19, SZL^{+}20, YZH^{+}21]. Compiler [SDD^{+}17]. Compiler-Inserted [SDD^{+}17]. Compilers [XLY^{+}22]. complete [CBvL24]. Completion [YX22]. Complex [GLH^{+}19]. Complexity [AGGS20, AGS17, WS17b, ZWT^{+}17]. Comprehensive [SATN22]. Compromise [GFS23]. Computation [DAR^{+}21]. Computationally [CT19]. Compute [JJS17]. Computer [LPF^{+}22]. Computers [SIR20]. Computing [KKT^{+}19, LGH^{+}17, RT24, TKZ^{+}18]. Concealing [SLWS20]. Conditions [SGHB20, ZTS18]. Configuration [KCMX20, ZCM^{+}22]. Configurations [SDA18]. Congested [WMX17]. Congestion [MSJ^{+}19, RGBY^{+}19, ZMSS19]. Congestion-Controlled [RGBY^{+}19]. Connected [ALBS^{+}20]. Connectivity [SAM^{+}22]. Conquer [LMSC22]. Consequences [VS19]. Consistency [LYQ^{+}22]. Constant [CMM22]. Constrained [CHT^{+}22, JLGJ^{+}24, KMM21, LCS^{+}21]. Constraint [HCZ^{+}21, TTL^{+}21]. Constraints [AN20, AGS17, LYP^{+}19, TG23, WYN20, WZS20, YHS^{+}20]. Consumption [BRC21, CNW^{+}17, YSL^{+}22]. Contemporary [JKR18]. Content [GKK20, SCH^{+}24, WM20]. Continuous [CPF20]. Contract [DAR^{+}21]. Contracts [CLS^{+}21, KMM21]. Contractual [WM20]. Control [AMSS19, AEKB20, CES19, DV18, FM22, KSZ23, LCS^{+}21, LYQ^{+}22, LF24, LGK22, MSJ^{+}19, WLdA^{+}17, ZMSS19, ZGB^{+}23]. Controlled [RGBY^{+}19]. Controller [GAES23]. Convection [KKRK22]. Convergence [DBS17, SCS22, WS17b, WYN18, ZQX^{+}23]. Converging [LF24]. Convex [AN20, SLJ19, YDH^{+}18, LGW20]. Cooperative [ALZ22, AGS17]. Coordinated [ZSPZ23]. Core [CKMP19]. Correlated [DCKG19]. Cost [AMA18, AI19, BNS17, BSA^{+}21, HC18, LJG^{+}22, PSLW22]. Cost-Aware [BSA^{+}21].

Cost-Efficient [AMA18, LJG^{+}22]. Costs [LYR22]. Counterfeit [GWX^{+}20]. Coupled [WYN18]. CPU [RVC^{+}19]. Criteria [TMM20]. Critically [HV21]. Cross [HLH23, OKB^{+}18, RJ23, RSJ24]. Cross-Layer [RJ23, RSJ24, OKB^{+}18]. Cross-Path [HLH23]. Crossbar [GTL^{+}17]. Crowdsourced [SVL20]. Crowdsourcing [SCYO21]. Cryptocurrency [FVB^{+}18, GWX^{+}20, TWFO20, WLZ^{+}22]. Cryptocurrency-themed [WLZ^{+}22]. Crystal [ASME19]. CSI [SDD^{+}17]. Curing [HC18]. Curtain [ASV^{+}21]. Curvature [SAM^{+}22]. Curvature-based [SAM^{+}22].

D [LKJK20, LGC^{+}18, HV18, ADGF^{+}24, KKRK22, LZGN22]. DaeMon [GHT^{+}23]. DAG [MYS^{+}22]. Dandelion [FVB^{+}18, VFV17]. Dangling [ZLL^{+}23]. Data [ARK^{+}22, AAJV20, AEKB20, ATN^{+}22, BSA^{+}21, CEF^{+}20, CAJK17, FS19, GHT^{+}23, GGYZ17, HKY^{+}19, HCZ^{+}21, IYRR18, JLGJ^{+}24, KKRK22, KLCS19, KSZ23, LJH^{+}23, LZC^{+}23, MLW^{+}23, PZB^{+}23, SIR20, TKZ^{+}18, TPK^{+}18, TKK21, WMMC23, ZKAV20, ZCW21, ZZCC17]. Data-Driven [ATN^{+}22]. Data-flow [LZC^{+}23]. Datacenter [AAS23, GZB^{+}21, ZGB^{+}23, ZLM^{+}18, ZL18]. Datacenter-scale [ZL18]. Datasets [HLP23]. DBMSes [SATN22]. DCN [WCX^{+}18]. DCNs [YCX^{+}19]. Deadline [TG23]. Debloating [KCMX20]. DECAF [IKS21]. December [AGU22, CCR21, CFR23]. Decentralized [XWG^{+}21, ZRC^{+}22]. Decision [WYN18]. Deconstructing [CNW^{+}17]. Deep [KKS22, LPF^{+}22, LJG^{+}22, MDS^{+}24, PZB^{+}23, WCX^{+}18, XLY^{+}22, YSL^{+}22]. Degradation [NT19]. Degree [ZWT^{+}18]. Delay [BZ18, LYE19, PSLW22, WS17b, WW20, ZWT^{+}17, ZWT^{+}18, ZTS18, Zub20]. Delay-Optimal [ZWT^{+}17, Zub20]. Delays [DBS17]. Demand [AS17, SSB18].

Democratizing [ITI^{+}24]. Demystify [ZYH^{+}19]. Demystifying [GLH^{+}19, WWH^{+}22]. Dense [ZSPZ23]. Density [LKJK20]. Departures [SYH^{+}22]. Dependency [KALB23]. Dependent [HBV19, ZHZ^{+}21]. Deployment [BSA^{+}19]. Depth [AMBK23, CMH^{+}17]. Descent [CSX17, SX19, WYN20]. Design [ADGF^{+}24, BNS17, CF23, GZB^{+}21, LKS^{+}17, TSLG^{+}20, ZCW21]. Design-Induced [LKS^{+}17]. Designing [CJC^{+}19, YHG^{+}17, ZWT^{+}17]. Desktop [SIR20]. Detailed [ANBK22, GYG^{+}18]. Detecting [CLS^{+}21, LZC^{+}23, SLWS20, XWG^{+}21, ZLL^{+}23]. Detection [BKO^{+}23, LPJ23a, WMMC23, WKWY19]. Deterministic [BNS17, WHBW22]. Device [ILL^{+}23, LYJ^{+}21]. Devices [BRC21, CYG^{+}17, KKS22]. Diagnosis [ZL18]. Diagnostic [MMS^{+}23]. Did [IBL^{+}19, BRC21]. Different [HLP23]. Differentially [Zho22]. Differentiated [ZLX^{+}23]. DiffForward [ZLX^{+}23]. Diffusing [YLX17]. Diffusions [EC18]. Dimensional [LML^{+}20, SX19, AMSS19, CBvL24]. Direct [WXW22]. Directed [LKE19]. Disaggregated [AZW^{+}23, GHT^{+}23]. Discipline [GV24]. Discrete [MSPI22]. Discretization [SBY19]. Disparities [ZHM^{+}21]. Dispatching [XGS24]. Disruptions [IMMP20]. Dissecting [IKS21]. Distributed [AGS17, CSX17, DBS17, FKP^{+}21, HNS19, JIAM19, LZJ23, LSX^{+}23, LVW19, MCS^{+}19, MDBvL17, RKTV22, SX19, SDA18, SSM^{+}17, XVK22, YSL^{+}22, KLCS19]. Distribution [NVGT20]. Distributions [CPR19]. Dive [MDS^{+}24]. Diverse [MPIZ17]. Diversity [OCW24]. Divide [LMSC22]. Divide-and-Conquer [LMSC22]. DL [PYL24]. DL-based [PYL24]. DM [ADGF^{+}24]. DNN [KKP^{+}22]. Does [NKY^{+}18, ZMC24].

Domains [ZLL^{+}23]. Don’t [CK23]. Drain [CMH^{+}17]. DRAM [CYG^{+}17, GKK20, GYG^{+}18, GLH^{+}19, LKS^{+}17]. DRAMs [KKRK22]. Dremel [ZCM^{+}22]. Driven [ASME19, ATN^{+}22, KKRK22, SIS17, WCX^{+}18]. Driving [ZHZ^{+}21]. Dropout [SCS22]. DSM [GKK20]. Dual [WYN20]. Duo [ZGB^{+}23]. during [LYL^{+}22]. Dynamic [FKP^{+}21, FZCL18, IMMP20, KCT18, LT22, LGK22, RBT21, SIN22, SQ20, TPK^{+}18]. Dynamical [LGK22]. Dynamics [FWB17, LF24, PPL^{+}21].

Each [KALB23]. Early [LGC^{+}18]. Eat [KB20]. ECI [AMA18]. ECI-Cache [AMA18]. Edge [AEKB20, KKS22, LGH^{+}17]. Editorial [AGU22, AGU23a, AGU23b, BD19, CCR21, CCR22a, CCR22b, CGZ17, CFR23, CFR24a, CFR24b, GKW21, TM20]. Editors [WA18]. Effective [LGH^{+}17, YWB19]. Efficiency [HLB^{+}23, LC23, SVL20]. Efficient [AMA18, BKO^{+}23, CT19, CCG21, GFL^{+}22, GHT^{+}23, GAES23, HNS19, LJG^{+}22, SCY23, YCX^{+}19, LSX^{+}23]. Elasticity [HLB^{+}23, MDBvL17]. Electric [SZL^{+}20]. Embedding [LZC^{+}23, HNS19]. Empirical [SATN22]. Enabling [CMH^{+}17, SIN22]. Encrypted [ASV^{+}21, BSA^{+}19]. End [BZY^{+}22, GWX^{+}20, JHWC21, TG23, WXW22]. End-to-End [JHWC21, TG23, BZY^{+}22, GWX^{+}20]. Endurance [AMA18]. Energy [AN20, CNW^{+}17, CMH^{+}17, LC23, RVC^{+}19]. Energy-Harvesting [AN20]. Enhanced [WSC^{+}24]. Enhancing [ALBS^{+}20, CAJK17, TKK21]. Enough [LYJ^{+}21]. Entanglement [NVGT20]. Enterprise [ARK^{+}22]. Enterprise-Grade [ARK^{+}22]. Entropic [LMS^{+}24]. Environment [AMBK23, LML^{+}20]. Environment-Aware [AMBK23]. Environments [NKY^{+}18]. EOSIO [HWW^{+}20]. Epidemic [HC19]. Epidemics [HC18]. Episodic [SBY19]. EPT [BKO^{+}23]. EPT-based [BKO^{+}23]. Equilibrium [WM20]. Era [WWH^{+}22]. Erdos [CKMP19, DCKG19]. Estimated [Gas17]. Estimation [AASS18, AS17, CT19, NNVD17, NKY^{+}18, PSS23, YNJS21]. Ethereum [CLS^{+}21]. Evaluation [BNS17, JKR18]. Evaluations [JCS^{+}24]. Event [CMH^{+}17]. Event-based [CMH^{+}17]. Everything [NBB^{+}19]. Everywhere [CWW^{+}23]. Evolving [LF24]. Exchange [XWG^{+}21]. Executing [AI19]. Exfiltration [SIR20]. Expected [Gas17]. Experience [BSA^{+}19, GGYZ17]. Experimental [CYG^{+}17, GYG^{+}18, GLH^{+}19]. Expert [LYR22, YCH^{+}22]. Expert-Calibrated [LYR22]. Explained [OKB^{+}18]. Explaining [ZHM^{+}21]. Exploiting [CAJK17, SIR20]. Exploration [OCW24]. Exposing [EC18]. Extensions [CP18, WLZ^{+}22]. Extra [YLX17]. Extreme [PLX22]. Extremely [SGHB20].

Failure [CJC^{+}19, ZLM^{+}18]. Fair [AIAR^{+}24, JIAM19, TP18]. Fairness [BCM17, IMMP20, SIN22, ZMC24]. False [DSR24]. Families [ZZM^{+}19]. Fast [BNS17, LML^{+}20, LJG^{+}22, LSX^{+}23, SQ20, WS17b]. Fast-Convergence [WS17b]. Fault [SDA18]. Fault-Tolerant [SDA18]. FCFS [GYSHB21]. Federated [CL24, ZZLL21]. FedQV [CL24]. Feedback [CES19, LZJ23, PSLW22, vdBBvL19]. Fi [LC23, VPK18]. Fiat [HLP23]. Fiedler [DE20]. Field [AG22, AG23, Gas17, GV17, HSV19, SSM^{+}17, Van19, Yin17, RBT21]. Filters [DSR24, SVL20]. Finding [MLHC20]. Fine [CMH^{+}17, NB23, XJ22]. Fine-Grained [XJ22, CMH^{+}17, NB23]. Finger [LZGN22]. Fingerprints [ALBS^{+}20]. Finite [CPR19, DBS17]. Finite-Buffer [CPR19]. Finite-Sum [DBS17]. First [KMM21, LC23, PLX22, Van22]. First-Come [Van22]. First-Served [Van22]. Fit [NKY^{+}18]. Flash [ARK^{+}22, CAJK17, LGC^{+}18]. Flash-based [CAJK17]. Flexibility [KKP^{+}22]. Flexible [LSX^{+}23, QTES19]. Flow [JIAM19, LZC^{+}23, QTES19]. Flows [MLW^{+}23]. Forecasting [FS19]. Fork [Zub20]. Fork-Join [Zub20]. Formal [FVB^{+}18]. Formalism [KKP^{+}22]. Forums [GFS23]. Forwarding [ZLX^{+}23]. Fragmentation [SSB18]. Framework [DMB^{+}22, LJG^{+}22, NB23, RJ23, SDD^{+}17, ZHZ^{+}21]. Free2Shard [RKTV22]. Fully [GHT^{+}23]. FuncPipe [LJG^{+}22]. Function [Zho22]. Fundamental [BBS20, FM19, WLS19]. Fusing [LYL^{+}22]. Fuzzing [HHP20].

G [CP18, SGHB20, SHBSW20, YS24].

Gacha [CF23]. Game

[BZY^{+}22, CF23, FZCL18, WS17a]. Game-Theoretic [FZCL18]. Gaming [IKS21]. Gap [CT19]. Garbage [ASME19]. Gazer [ASME19]. GB [BRC21]. General [SGHB20]. Generalized [ZZM^{+}19]. Generation [JLGJ^{+}24]. Geo [KLCS19]. Geo-distributed [KLCS19]. Geographic [AGS17]. GHz [ANBK22]. GI [CP18]. GI/G/1 [CP18]. Gittins [SGHB20]. Given [YX22]. Global [JHWC21, Van19, ZQX^{+}23]. Go [GAES23, BRC21]. Go-to-Controller [GAES23]. Gossiping [ZZLL21]. Got [AZG24]. GPGPU [YNJS21]. GPU [JKR18, KKRK22, SATN22]. GPU-Driven [KKRK22]. GPUs [ADGF^{+}24, DMB^{+}22, LPJ23b, TPK^{+}18]. Grade [ARK^{+}22]. Gradient [CSX17, DBS17, SX19, WLS19]. Grained [XJ22, CMH^{+}17, NB23]. Graph [CKMP19, EC18, JJS17]. Graphs [DCKG19, HC19, LKE19, SZL19, VSS22, YXL21]. Great [MSJ^{+}19, ZMW^{+}20]. Greedy [TTL^{+}21]. Grid [OCW24].

Ground [HHP20]. Ground-Truth [HHP20]. Growth [SVL20]. GuaNary [BKO^{+}23]. Guarantee [ZCW21]. Guarantees [FVB^{+}18, KCT18, LSNI21, MBLYR23]. Guardrails [GSHB19].

H [ADGF^{+}24]. Hadoop [GGYZ17]. Happen [DSR24]. Hardware [BG23, GKK20, LYJ^{+}21]. Hardware-Assisted [GKK20]. Hardware-Aware [LYJ^{+}21]. Harvesting [AN20, LZC^{+}23]. Hashing [SQ20]. haul [CWW^{+}23, KJM^{+}22]. HEAL [PZB^{+}23]. Heart [KCMX20]. Heavy [GSHB19, GSHBSW22, SvKB^{+}20, XGS24, Yin17, ZWT^{+}17, ZWT^{+}18, ZTS18]. Heavy-Tailed [SvKB^{+}20]. Heavy-Traffic [XGS24, ZWT^{+}17, ZWT^{+}18, ZTS18]. Hegemony [ZHM^{+}21]. Heterogeneous [AG22, AAJV20, ZKAV20]. Hierarchical [YCH^{+}22]. Hieroglyph [JJS17]. High [AMA18, ADGF^{+}24, HKY^{+}19, LPF^{+}22, LKJK20, SX19, ZGB^{+}23]. High-bandwidth [ADGF^{+}24]. High-capacity [ADGF^{+}24]. High-Density [LKJK20]. High-Endurance [AMA18]. High-Performance [HKY^{+}19, LPF^{+}22]. High-Throughput [ZGB^{+}23]. Homophilous [ZHM^{+}21]. Hop [CWW^{+}23]. hops [YXL21]. Horizon [SCY23]. Host [WXW22]. Hosting [ZLL^{+}23]. Hosting-Based [ZLL^{+}23]. Hosts [PZB^{+}23]. Hound [ZL18]. Hybrid [ADGF^{+}24, ASME19]. Hyper [vdBBvL19]. Hyper-Scalable [vdBBvL19]. HyperBench [WZT19]. Hyperexponential [Van19].

I/O [AMA18, RT24, WXW22]. iBox [ATN^{+}22]. Identify [ZHZ^{+}21]. If [PSC19]. Illuminating [HLP23]. Images [AMBK23]. Imbalance [PJDQ20, ZWT^{+}18]. Impact [GV24, PST17, YHG^{+}17]. Implications [AGGS20]. Impression [KMM21]. Improvement [Van22]. Improving [AAJV20, GYSHB21, LGC^{+}18]. In-depth [CMH^{+}17]. Increasingly [ASV^{+}21]. Index [AdV18, LYL^{+}22]. Induced [LKS^{+}17]. Industrial [WLdA^{+}17]. Inequalities [ABG^{+}20]. Inference [Cas17, ILL^{+}23, KSZ23, SMF^{+}22]. Inferring [BSA^{+}19, SK19]. Influence [DDT^{+}22, NNVD17]. Information [AMBK23, BZ18, LCS^{+}21, YX22]. Infrastructure [CWW^{+}23]. Infrastructures [SAM^{+}22]. Input [FKP^{+}21, GTL^{+}17, YNJS21]. Input-Dynamic [FKP^{+}21]. Input-Queued [GTL^{+}17]. Insensitivity [KV21]. Inserted [SDD^{+}17]. Instances [WUNK17]. Instrumentation [SDD^{+}17]. Intel [BKO^{+}23, NBB^{+}19, OKB^{+}18]. Intelligent [JP23]. Intensive [LJH^{+}23]. Interacting [DE20]. Interactions [GLH^{+}19]. Interactive [AI19, LZGN22]. Intercontinental [CWW^{+}23]. Internet [MSJ^{+}19, AZG24, PJDQ20, RSJ24, SK19, ZMW^{+}20]. Internet-wide [AZG24]. Interrupts [CK23]. Intersection [ZZCC17]. Interval [IBL^{+}19]. Inventories [LMSC22]. Inventory [LYP^{+}19, YHS^{+}20]. Inverse [FS20]. Invertible [MBLYR23]. Investigation [LML^{+}20, QWQ17]. IoT [LZGN22]. IP [RJ23]. IPFS [SCH^{+}24]. IPv6 [HLP23]. Issues [TL20]. Item [HXW23]. Iteration [ZQX^{+}23].

JavaScript [NB23]. Jetson [KKS22]. Job [SvKB^{+}20, Van19, GSHBSW22]. Jobs [WXHB21, WW20]. Join [HSV19, Zub20]. Join-Below-Threshold [HSV19]. Joint [FM22]. JS [NB23]. JSQ [vdBBvL19]. June [AGU23b, CCR22b, CFR24b].

Keeping [GSHB19]. Kernel

[CK23, KCMX20]. Kernelized [LZJ23]. Key [HKY^{+}19, SQ20]. Key-value [SQ20].

Kingman [CP18]. Knapsack [HCZ^{+}21, SZL^{+}20, SYH^{+}22, TTL^{+}21, YZH^{+}21]. Know [FS20, IBL^{+}19, NBB^{+}19]. KV [ZCM^{+}22]. KV-Store [ZCM^{+}22].

Labeling [DCKG19]. Labs [MMS^{+}23]. Lancet [CJC^{+}19]. Large [CLE^{+}24, CNK^{+}23, HV21, HXW23, LML^{+}20, MDBvL17, SSM^{+}17]. Large-Scale [HV21, LML^{+}20, MDBvL17]. Last [IBL^{+}19]. Latency [LKS^{+}17, LC23, PJDQ20, TKFJ23]. Latent [SCY23]. Law [YXL21]. Layer [RJ23, RSJ24, OKB^{+}18]. Learning [ALZ22, BSA^{+}21, CES19, CSX17, CL24, FPCS22, FM22, HC19, KKS22, LYR22, LPF^{+}22, LJG^{+}22, LF24, MPIZ17, RS22, SCY23, SGS19, SBY19, SZL19, SX19, TP18, WCX^{+}18, WYN18, XLY^{+}22, YDH^{+}18, YCH^{+}22, YSL^{+}22, ZQX^{+}23, ZAA^{+}24, ZL18, Zho22]. Least [HV18]. Leave [ALBS^{+}20]. Length [KCT18, LM18]. Length-Bounded [KCT18]. LEO [ITI^{+}24]. Lessons [GYG^{+}18]. Level [CK23, KB20]. Leveraging [CL24, HLB^{+}23, LZGN22]. Lifetime [CAJK17, LGC^{+}18]. Light [Yin17, YS24]. Light-Tailed [YS24]. Lightened [JP23]. Lightweight [FVB^{+}18, MLW^{+}23]. Like [IYRR18]. Limitation [ZWT^{+}18]. Limitations [JCS^{+}24]. Limits [BBS20, FM19, WLS19]. Linear [FS20, LYQ^{+}22, LGK22, SCS22, YHS^{+}20, YX22, Zho22]. Links [RJ23]. Linux [QWQ17]. Listing [MBLYR23]. Live [SLWS20]. Load [ALZ22, ALR17, CNW^{+}17, GSHB19, HBV19, HV21, HSV19, LCZ^{+}23, MCS^{+}19, MBvL18, PJDQ20, PSC19, TS19, WZS20, YLX17, ZWT^{+}17, ZWT^{+}18, ZTS18]. Load-Balanced [PJDQ20, YLX17]. Loaded [HV18, HV21]. Loading [LYL^{+}22]. Loads [YLX17]. Local [SZL19, ZGB^{+}23]. Locality [TPK^{+}18, TKK21, WZS20]. Localized [ZQX^{+}23]. Locally [JJS17].

Locally-Sufficient [JJS17]. Location [LZC^{+}23, PTD20, SMF^{+}22, YWB19]. Location-Based [YWB19]. Logarithmic [LVW19]. Long [CWW^{+}23, SIN22, SCY23]. Long-haul [CWW^{+}23]. Long-term [SIN22]. Longevity [CAJK17]. Look [ANBK22, ASV^{+}21, LC23, SLJ19, SCH^{+}24]. Look-Ahead [SLJ19]. Lookup [LSX^{+}23, MBLYR23]. Lookups [SQ20]. Loss [CPR19, LGC^{+}18]. Losses [LGW20]. Loud [IYRR18]. Low [AGS17, GSHB19, LYE19, LKJK20, SCY23, TP18, WS17b, ZWT^{+}17]. Low-Complexity [AGS17, WS17b, ZWT^{+}17]. Low-Delay [WS17b]. Low-Rank [SCY23]. Low-Wear [LKJK20]. LPM [GAES23]. LRU [QTES19, TQJS18]. LTE [PLX22, TLL^{+}18, VPK18]. LTE-Aware [VPK18]. LTERadar [VPK18]. Ludo [SQ20]. Lux [HLP23].

M [WHBW22, SGHB20, SHBSW20, WHBW22, YS24]. M/G/ [SGHB20]. M/G/1 [SHBSW20, YS24]. M/M/ [WHBW22]. Machine [BSA^{+}21, CSX17, ZAA^{+}24]. Magma [HHP20]. Malcolm [ALZ22]. Malicious [WLZ^{+}22]. MalRadar [WWH^{+}22]. Malware [CNK^{+}23, WWH^{+}22]. Management [ALZ22, LNY21, QTES19, SIS17, YWH17, YHS^{+}20]. Many [BZ18]. Many-Sources [BZ18]. Mapping [MDS^{+}24]. March [CCR22a, AGU23a, CFR24a]. Marketplace [AZW^{+}23]. Marketplaces [ABG^{+}19]. Markov [CT19, DV18, WYN18, WKWY19]. Markov-Modulated [DV18]. Markovian [LKE19]. Mars [AAS23]. Match [ABG^{+}20, TKK21]. Matching [YXL21]. Matrix [AASS18, GFL^{+}22, MCS^{+}19]. Max [JIAM19]. Max-Min [JIAM19]. Maximization [CHT^{+}22, TTL^{+}21]. Maximize [FPCS22]. Maximizing [DAR^{+}21]. Mean [AG22, AG23, Gas17, GV17, HV21, HSV19, RBT21, SSM^{+}17, Van19, Yin17]. Mean-Field [Gas17, SSM^{+}17, RBT21]. Measurement [BRC21, HHM^{+}23, ITI^{+}24, LPJ23a, MMS^{+}23, ZTY^{+}21, ZZCC17, ZZM^{+}19]. Measurements [JHWC21, NB23]. Measuring [LZC^{+}23, LTS23, PPL^{+}21, PTD20, ZLL^{+}23, AZG24]. Mechanism [TSLG^{+}20]. Mechanisms [CEF^{+}20, CYG^{+}17, LKS^{+}17]. Meets [WCX^{+}18]. Membership [PTD20]. Memories [ASME19]. Memory [ADGF^{+}24, AZW^{+}23, CAJK17, GFL^{+}22, JCS^{+}24, LGC^{+}18, NB23, RKK22, SZL19, ZKAV20]. Memtrade [AZW^{+}23]. Merge [JJS17]. Merging [GKK20]. Message [WA18]. Metamorphic [XLY^{+}22]. Metaverse [LTS23]. MetaVRadar [LTS23]. Method [ALR17, Yin17]. Methodology [JKR18]. Methods [Cas17, DBS17]. Metric [SBY19]. Metrics [NT19]. Microarchitecture [PYL24]. Middle [ZYH^{+}19]. Millimeter [RS22]. Millimeter-Wave [RS22]. MIMO [ANBK22]. Min [JIAM19]. Mini [ZTY^{+}21]. Mini-Apps [ZTY^{+}21]. Minimal [IMMP20]. Minimization [LGK22]. Minimizing [LM18, QTES19]. Miracle [HHM^{+}23]. Mirage [HHM^{+}23]. Mirror [WYN20]. Mis [HWW^{+}20]. Misinformation [WKWY19]. Misses [QTES19]. mitigate [ZMC24]. Mitigating [RT24]. Mix [TKK21]. Mixed [AI19]. ML [KSZ23]. mmWave [LZGN22, ZSPZ23]. Mobile [BZY^{+}22, BRC21, CNW^{+}17, LZC^{+}23, NB23, TLL^{+}18]. Mobility [PLX22, SSB18]. Mobility-on-Demand [SSB18]. Model [AASS18, CMH^{+}17, CLE^{+}24, GSHBSW22, ILL^{+}23, KV21, LGH^{+}17, PST17, SSM^{+}17, ZRC^{+}22]. Modeling [RVC^{+}19, WLdA^{+}17]. Models [AG23, BSA^{+}19, GYG^{+}18, KKS22, LM18, LJG^{+}22, Van19]. Modem [CMH^{+}17]. Modern [BG23, CYG^{+}17, LKS^{+}17, ZLX^{+}23]. Modified [TTL^{+}21]. Modulated [DV18]. Monetizing [XVK22]. Monitoring [CV23, IBL^{+}19]. Monotone [TTL^{+}21]. Monte [LKE19]. Motion [LZGN22]. Movement [GHT^{+}23]. MPX [OKB^{+}18]. mRSC [AMSS19]. Multi [ALZ22, AMSS19, CBvL24, CV23, IYRR18, ILL^{+}23, LVW19, QTES19, SGS19, TS19, VSS22, WXHB21, YCH^{+}22, YPA^{+}18, ZQX^{+}23]. Multi-Agent [LVW19, VSS22, ZQX^{+}23, ALZ22]. Multi-armed [YPA^{+}18]. Multi-dimensional [AMSS19, CBvL24]. Multi-flow [QTES19]. Multi-Model [ILL^{+}23]. Multi-Pipe [CV23]. Multi-scale [YCH^{+}22]. Multi-Server [TS19, WXHB21]. Multi-tenant [IYRR18]. Multicut [KCT18]. Multidimensional [YZH^{+}21]. Multihop [TG23]. Multilayer [JNTG18]. Multipartite [NVGT20]. Multiple [LMSC22, SZL^{+}20, WS17a]. Multiplexing [BNS22]. Multiplication [GFL^{+}22, MCS^{+}19]. Multiscale [AG23]. Multiserver [GSHBSW22]. Multiserver-job [GSHBSW22]. My [AZG24, SIR20, BRC21].

N1 [CCR22a, AGU23a, CFR24a]. N2 [AGU23b, CCR22b, CFR24b]. N3 [AGU22, CCR21, CFR23]. Named [GGYZ17]. NAND [LKJK20, LGC^{+}18]. Nautilus [RJ23]. NDT7 [MMS^{+}23]. Near [AAS23, HXW23, MCS^{+}19, SHBSW20, TKZ^{+}18, TG23, WSC^{+}24]. Near-Optimal [AAS23, HXW23, SHBSW20, TG23, WSC^{+}24]. Near-Perfect [MCS^{+}19]. Nearly [SGHB20, YX22]. Necessary [ZTS18]. Negative [DSR24]. NetDiffusion [JLGJ^{+}24]. Network [ATN^{+}22, BBS20, BNS17, BNS22, BSA^{+}21, CJC^{+}19, CV23, DDT^{+}22, FWB17, FM19, IBL^{+}19, ITI^{+}24, JLGJ^{+}24, KJM^{+}22, KCT18, LM18, LPJ23a, LTS23, MMS^{+}23, NGB^{+}21, NT19, SAM^{+}22, SQ20, STGM21, VFV17, WS17b, WCX^{+}18, YHG^{+}17, ZGB^{+}23, ZMC24, ZZCC17, ZZM^{+}19]. Network-Caching [BBS20]. Networked [YHW23, ZQX^{+}23]. Networking [CK23, FVB^{+}18, GGYZ17, WXW22, ZGN^{+}19]. Networks [AAS23, ALR17, AdV18, BZ18, FKP^{+}21, FM22, GBvL24, JNTG18, LSNI21, NNVD17, OCW24, RGBY^{+}19, SCS22, TLL^{+}18, TWFO20, TKFJ23, TL20, TG23, WMX17, XJ22, ZLM^{+}18, ZHM^{+}21, ZSPZ23]. Neural [LYJ^{+}21, NGB^{+}21, SCS22, WCX^{+}18, WSC^{+}24]. Neural-Enhanced [WSC^{+}24]. NextG [XJ22]. NFT [HHM^{+}23]. NG [XJ22]. NG-Scope [XJ22]. No [MSPI22, ALBS^{+}20]. No-regret [MSPI22]. Node [PST17]. Noise [DDT^{+}22, SIR20]. Noisy [HC19]. Non [CBvL24, DMB^{+}22, LKE19, LGW20, LGK22, PG17, WS17a, YDH^{+}18]. Non-additive [WS17a]. Non-complete [CBvL24]. Non-Convex [YDH^{+}18, LGW20]. Non-Markovian [LKE19]. Non-Preemptive [PG17]. Non-stationary [LGK22]. Non-Uniform [DMB^{+}22]. Nonideal [TL20]. Nonlinear [PSLW22]. Nonlinearity [EC18]. Novel [LKJK20]. NTP [MDS^{+}24]. Nudge [Van22, GYSHB21]. NURA [DMB^{+}22].

O [AMA18, AIAR^{+}24, RT24, WXW22]. O-RAN [AIAR^{+}24]. Obfuscating [ALBS^{+}20]. Object [BBHB18]. Observations [YPA^{+}18]. ODE [Van19]. ODE-based [Van19]. Offline [LNY21]. On-Device [ILL^{+}23]. Once [BNS22]. One [LYJ^{+}21, NKY^{+}18, SHBSW18]. Online [ABG^{+}19, AN20, BBS20, CYC^{+}19, LNY21, LCZ^{+}23, LCS^{+}21, LSNI21, LYR22, LYP^{+}19, LGW20, LMSC22, MPIZ17, PSLW22, RCMW23, SLJ19, SZL^{+}20, SYH^{+}22, TSLG^{+}20, TMM20, WYN18, WYN20, YWH17, YDH^{+}18, YHS^{+}20, YZH^{+}21, YLIR24, YHW23, ZLW17]. Only [BNS22].

Ookla [MMS^{+}23]. Open [ARK^{+}22]. Open-Source [ARK^{+}22]. Operating [KCMX20]. Operation [CYG^{+}17]. Optical [ZCW21]. Optimal [AAS23, BBHB18, DV18, GAES23, GSHBSW22, HXW23, LCZ^{+}23, LF24, LMS^{+}24, MDBvL17, MBvL18, SvKB^{+}20, SGHB20, SHBSW20, TMM20, TG23, WMX17, WSC^{+}24, WKWY19, WZS20, XGS24, YWH17, YDH^{+}18, YS24, ZLW17, ZKAV20, ZWT^{+}17, CPF20, Zub20]. Optimality [BDHB17, ZWT^{+}18, ZTS18]. Optimally [MPIZ17]. Optimistic [MSPI22]. Optimization [AdV18, AN20, CYC^{+}19, DBS17, LYR22, LYP^{+}19, LGW20, LMSC22, LVW19, PSLW22, RCMW23, SLJ19, WS17b, YHS^{+}20]. Optimized [HLH23]. Optimizing [AI19, HLB^{+}23]. Option [SZL19]. Order [CPF20]. Order-optimal [CPF20]. Outbound [MLHC20]. Outward [NNVD17]. Overcoming [SCY23, ZWT^{+}18]. Overflow [BKO^{+}23]. Overlapping [TMM20]. Own [KALB23].

Pace [KALB23]. Packet [NPAP19, TG23]. Packing [HXW23, LT22]. Page [BKO^{+}23, CNW^{+}17, LYL^{+}22, PSC19]. PageRank [VS19]. Paradigm [QTES19]. Parallel [BDHB17, LGH^{+}17, WW20]. Parallelism [TPK^{+}18]. Partial [CKMP19, Zub20]. Partition [LGH^{+}17]. Party [CEF^{+}20, KALB23, PBLC21]. Patching [WLdA^{+}17]. Path [ATN^{+}22, HLH23, KSZ23, LZC^{+}23, STGM21]. Paths [PJDQ20]. Pause [LCZ^{+}23]. Paying [BNS22]. Payment [TWFO20]. PDE [ALR17]. PEACH [AMBK23]. Peer [TKFJ23]. Peer-to-Peer [TKFJ23]. Peering [WM20]. Per-Flow [JIAM19]. Perfect [MCS^{+}19]. Performance [ANBK22, AAJV20, BCM17, Cas17, DDT^{+}22, HKY^{+}19, HBV19, IKS21, KKS22, LPF^{+}22, LPJ23a, NBB^{+}19, PZB^{+}23, PBLC21, RVC^{+}19, SCH^{+}24, SATN22, WXW22, ZCW21, ZGN^{+}19, ZMW^{+}20]. Persistent [ZZCC17]. Personalized [VS19]. Perspective [GZB^{+}21]. Phishing [CNK^{+}23]. Physical [ZCW21]. Picocells [RS22]. Pipe [CV23]. Pipelined [LJG^{+}22]. Placement [KKRK22, LSX^{+}23, PST17, ZKAV20]. Plane [MLW^{+}23]. Platforms [AMA18, AIAR^{+}24, BZY^{+}22]. Point [LPJ23a]. Points [VPK18]. Policies [HBV19, SHBSW18, SvKB^{+}20, Zub20]. Policy [SGHB20, ZQX^{+}23]. Polynomial [CCG21]. Polynomial-Time [CCG21]. POMACS [AGU22, AGU23a, AGU23b, CCR21, CCR22a, CCR22b, CFR23, CFR24a, CFR24b]. Ponzi [CLS^{+}21]. Pool [MDS^{+}24]. Pooling [CBvL24, TQJS18]. Pools [BCM17]. Popularity [MDS^{+}24]. Portfolio [SIS17]. Portfolio-driven [SIS17]. Posted [ZLW17]. Potential [ZRC^{+}22]. PoW [DAR^{+}21]. Power [BNS22, CMH^{+}17, EC18, FWB17, GYG^{+}18, GZB^{+}21, HV18, HV21, IYRR18, OCW24, SCYO21, YXL21]. Power-Law [YXL21]. Power-of-d-choices [HV18]. Practical [BBHB18, BSA^{+}19, SQ20]. Practicality [KCMX20]. Practices [JHWC21, WLdA^{+}17]. Predict [ABG^{+}20]. Predictable [RS22]. Prediction [AMBK23, NGB^{+}21, STGM21, YSL^{+}22, ZLM^{+}18]. Predictions [CYC^{+}19, LYQ^{+}22, LGW20, LYE19, LT22, RCMW23]. Predictive [KJM^{+}22, YHG^{+}17]. PredictRoute [STGM21]. Preemptive [PG17]. PreFix [ZLM^{+}18]. Price [KMM21, PPL^{+}21, SSB18]. Prices [ZLW17]. Pricing [SCYO21, WMX17]. Primal [WYN20]. Primal-Dual [WYN20]. Privacy [ALBS^{+}20, PPL^{+}21, PTD20, TWFO20]. Privacy-Utility [TWFO20]. Private [SAM^{+}22, WM20, Zho22, LZJ23]. Proactive [AMBK23, LYE19]. Problem [FS20, KCT18, LCZ^{+}23, SZL^{+}20, SYH^{+}22].

Problems [YZH^{+}21, YCH^{+}22]. Process [LJH^{+}23, LGC^{+}18, PSS23]. Processes [CES19, WYN18]. Processing [GFL^{+}22, JCS^{+}24, JJS17, LPJ23b]. Processing-In-Memory [GFL^{+}22, JCS^{+}24]. Processor [KV21]. PROFET [RVC^{+}19]. Profile [ASME19]. Profile-Driven [ASME19]. Profiling [PSC19]. Program [SDD^{+}17]. Properties [LZGN22]. Prophet [ABG^{+}20]. Proportional [FZCL18, GTL^{+}17]. Proportionally [TP18]. Protection [BKO^{+}23, GFS23]. Protocol [JLGJ^{+}24]. Protocol-Constrained [JLGJ^{+}24]. Providers [CEF^{+}20, WM20, YWB19]. Provisioning [CYC^{+}19]. Proxy [LYJ^{+}21]. Pruned [CJC^{+}19]. Pseudonymous [GFS23]. Public [WUNK17]. Pull [ZTS18]. Pull-based [ZTS18]. Pulls [HHM^{+}23].

QoE [AdV18]. QoS [NT19, YWH17]. Quadratic [CL24, LYQ^{+}22]. Quality [BNS17, BSA^{+}19]. Quantifying [TPK^{+}18]. Quantitative [JKR18]. Queries [IBL^{+}19]. Query [SATN22]. Queue [BZ18, CPR19, GTL^{+}17, LM18, ZWT^{+}18]. Queue-Proportional [GTL^{+}17]. Queued [GTL^{+}17]. Queueing [FM22, WXHB21, WW20]. Queues [CPR19]. Quickest [WKWY19]. QuickStop [WKWY19].

Rack [ALZ22]. Radiometric [ALBS^{+}20]. Rain [KJM^{+}22]. RAN [AIAR^{+}24]. Random [DE20, TS19, Zub20]. Randomized [ALR17, YWH17, YLX17]. Ranging [MLW^{+}23]. Rank [SCY23]. Rankings [MPIZ17]. Rate [DBS17, SCS22]. Rateless [MCS^{+}19]. Rates [DSR24, JIAM19]. Rationing [ASME19]. Re [HNS19, PYL24]. Re- [HNS19]. Re-Thinking [PYL24]. Reaching [KCT18]. Real [GFL^{+}22, JCS^{+}24, KMM21, PHSB21, WMMC23]. Real-time [KMM21, PHSB21, WMMC23]. Reality [LTS23]. Receipts [NPAP19]. Reconfigurable [AAS23, YCX^{+}19, ZGB^{+}23]. Reconfiguration [KJM^{+}22]. Reconnaissance [HLH23]. Recovery [CKMP19]. Recursive [BCM17]. Recycling [DSR24, RKK22]. Redesigning [VFV17]. Reduced [CYG^{+}17]. Reduced-Voltage [CYG^{+}17]. Reduction [ARK^{+}22, LKS^{+}17, TKFJ23]. Redundancy [AAJV20, GV24, RB20, Zub20]. Refined [AG22, GV17]. Refinement [AG23]. Refinery [BSA^{+}21]. Refining [RBT21]. Regimes [Yin17]. Register [ZZCC17]. Regret [BBS20, CPF20, LM18, LGK22, TP18, MSPI22]. Regrets [CYC^{+}19]. Regulators [TL20]. Reinforcement [SCY23, SBY19, ZQX^{+}23, Zho22]. Reliable [LKJK20]. Remote [PSS23]. Renewal [CES19]. Rényi [CKMP19, DCKG19, LMS^{+}24]. Reorganizing [TKK21]. Replenishable [YLIR24]. Replication [SSM^{+}17]. Reports [CNK^{+}23]. Representation [BSA^{+}21]. Reproducing [ZHM^{+}21]. Resilience [CJC^{+}19, KCT18, NT19, RSJ24, YNJS21]. Resilient [YHG^{+}17]. Resistances [YWB19]. resistant [RKTV22]. Resource [AIAR^{+}24, BCM17, CBvL24, CYC^{+}19, DMB^{+}22, FPCS22, HSV19, LSX^{+}23, RKTV22, SIN22, SIS17, TQJS18, TSLG^{+}20, YSL^{+}22, ZLW17]. Resource-efficient [LSX^{+}23]. Resources [WS17a]. Response [GSHB19, SvKB^{+}20]. Restless [DV18]. Result [VS19]. Results [GGYZ17]. Resume [LCZ^{+}23]. Retail [AS17]. Retention [LGC^{+}18]. Retroactive [NPAP19]. Reusable [FPCS22]. Revisiting [TTL^{+}21]. Richer [NGB^{+}21]. Risks [ZLL^{+}23]. Road [GSHB19]. Robotics [BG23]. Robust [AMSS19, VSS22]. Robustness [LYQ^{+}22]. RocksDB [ZCM^{+}22]. Root [LML^{+}20]. Router [SDA18]. Routing [PHSB21, PST17, SK19, TWFO20, YCX^{+}19, ZGB^{+}23]. Rows [GKK20]. RPC [KSZ23]. RTBH [AZG24]. Rug [HHM^{+}23]. Ruin [CEF^{+}20]. Rules [ZHZ^{+}21].

SADPonzi [CLS^{+}21]. Safe [YLX17]. Same [GKK20]. Sample [SCY23]. Sample-Efficient [SCY23]. Samples [YX22]. Sampling [GTL^{+}17, NPAP19, PSS23, TS19]. Satellite [ITI^{+}24]. Saturation [GBvL24]. SCADA [OCW24]. Scalability [DDT^{+}22, GSWV20, JCS^{+}24]. Scalable [vdBBvL19]. Scale [ALZ22, CLE^{+}24, CNK^{+}23, HV21, KCT18, LML^{+}20, MDBvL17, QWQ17, NNVD17, YCH^{+}22, ZL18]. Scaling [BZ18]. Scam [LGL23, XWG^{+}21]. Scenario [ZHZ^{+}21]. Scenario-dependent [ZHZ^{+}21]. Scenarios [CBvL24]. Schedules [LPJ23a]. Scheduling [BDHB17, GTL^{+}17, GSHBSW22, JP23, PG17, SHBSW18, SHBSW20, TG23, Van22, YS24]. Scheme [AMA18]. Schemes [CLS^{+}21, ZWT^{+}17]. Scientific [RT24]. Scope [XJ22]. Search [LYJ^{+}21]. Second [KMM21]. Securing [SX19]. Security [ALBS^{+}20, LC23, WS17a, ZYH^{+}19, ZLL^{+}23]. Segmentation [ABG^{+}19]. Segmentation-Thickness [ABG^{+}19]. Selection [HV18]. Semi [CCG21]. Semi-Bandits [CCG21]. Sensitive [TL20]. Separation [TQJS18]. Series [AASS18, PTD20]. Served [Van22]. Server [GBvL24, HV18, TS19, WXHB21]. Serverless [LJG^{+}22, MYS^{+}22, RT24]. Servers [HSV19, SIS17]. Service [AZG24, GV24, HXW23, LML^{+}20, MDBvL17, YWB19]. Services [LYE19, SSB18, ZLX^{+}23]. Serving [KSZ23]. Set [KCMX20, NKY^{+}18]. Sets [CJC^{+}19]. Settings [CSX17]. Setup [WHBW22]. SGD [CMM22]. SGX [NBB^{+}19]. Shallow [AAS23, SCS22]. Sharing [FZCL18, HSV19, KV21, ZRC^{+}22]. Shifting [LCZ^{+}23]. Should [NBB^{+}19]. Shrinking [LMS^{+}24]. Shuffle [LJH^{+}23]. Side [IYRR18, YX22]. Sided [WMX17]. Signal [SIR20]. Signals [LZGN22]. SimNet [LPF^{+}22]. Simple [CEF^{+}20, LGH^{+}17, SHBSW20]. Simulating [RVC^{+}19]. Simulation [ATN^{+}22, JKR18, LPF^{+}22, PYL24]. Size [NNVD17, NKY^{+}18, XGS24]. Size- [XGS24]. Sizes [BBHB18, HXW23, SvKB^{+}20, Van19]. Sizing [YNJS21]. Sketch [ZZM^{+}19]. Skewed [GBvL24]. SLITS [JP23]. Slowdowns [Zub20]. Smart [BRC21, CLS^{+}21, DAR^{+}21, PHSB21]. SmartNICs [LJH^{+}23]. Smash [LSX^{+}23]. Smoothed [RCMW23]. Smoothness [CPF20]. SOAP [SHBSW18]. Social [SGS19, ZMC24]. Software [ZHZ^{+}21]. Solving [FS20]. Some [IYRR18]. Source [ARK^{+}22]. Sources [BZ18]. Space [CBvL24, RKK22]. Spaces [SBY19]. Spare [XVK22]. Sparse [GFL^{+}22, vdBBvL19]. SparseP [GFL^{+}22]. Sparsity [JP23]. Sparsity-Lightened [JP23]. Spectral [CT19]. Speed [LYL^{+}22]. Speeding [YNJS21]. Speedtest [MMS^{+}23]. Splicing [GAES23]. SplitRPC [KSZ23]. Splitting [KSZ23]. Spread [WMMC23, ZZCC17]. SSD [LNY21]. Stability [GSWV20, RB20]. Stabilization [YHW23]. Stack [CK23, KSZ23, OKB^{+}18, QWQ17]. Staleness [AEKB20]. StarShip [RT24]. State [AMBK23, BZ18, CBvL24, JIAM19, RBT21, XGS24]. State-Aware [XGS24]. Stationary [CMM22, LGK22]. Statistical [CSX17, SX19]. Statistically [CCG21]. Stay [ALBS^{+}20]. Stealthy [SIR20]. Stein [Yin17]. Stepsize [CMM22]. Stochastic [BNS22, FM22, HXW23, JNTG18, SSM^{+}17, Van22, WYN20]. Stochastically [GYSHB21]. Stopping [WKWY19]. Storage [ARK^{+}22, BRC21, CAJK17, LKJK20, LSX^{+}23, SSM^{+}17, ZKAV20]. Store [HKY^{+}19, ZCM^{+}22]. Straggler [ZL18]. Strategic [TKFJ23]. Strategies [PST17, TMM20]. Streaming [BZY^{+}22, BSA^{+}19, CHT^{+}22, KLCS19, NGB^{+}21, WMMC23, WSC^{+}24]. Strongly [YS24]. Structural [VS19]. Structure [RGBY^{+}19, SCY23]. Structures [HKY^{+}19]. Study [BG23, CNK^{+}23, FWB17, GV24, GYG^{+}18, GLH^{+}19, HHM^{+}23, KKRK22, PLX22, SATN22, WYN18, ZTY^{+}21]. style [AN20]. Sub [BKO^{+}23]. Sub-Page [BKO^{+}23]. Submarine [RJ23]. Submodular [CHT^{+}22, HCZ^{+}21, TTL^{+}21]. Substation [FWB17]. Sufficient [JJS17, ZTS18]. SUGAR [YNJS21]. Suite [WZT19]. Sum [DBS17]. Summarization [HCZ^{+}21]. Summer [IBL^{+}19]. Supermarket [KV21]. Supply [ABG^{+}20]. Support [BKO^{+}23, GHT^{+}23]. Supporting [DMB^{+}22, TLL^{+}18]. Surface [SLWS20]. Switch [NVGT20, ZLM^{+}18]. Switched [ZCW21]. Switches [CV23, GTL^{+}17]. Switching [KJM^{+}22, LYR22, PSLW22, SIR20, YLX17]. Sync [JJS17]. Synchronization [TL20]. Synthesis [SDA18]. Synthetic [AMSS19]. System [JCS^{+}24, KCMX20, OKB^{+}18, RVC^{+}19]. Systematic [BG23, ZHZ^{+}21]. Systems [ARK^{+}22, AG22, Cas17, GV24, GHT^{+}23, GSWV20, HV21, HXW23, KV21, LF24, LVW19, LGK22, MDBvL17, NBB^{+}19, PHSB21, RB20, SQ20, SSM^{+}17, TS19, WLdA^{+}17, YSL^{+}22, YHW23, ZKAV20, ZAA^{+}24, ZWT^{+}18, ZTS18, Zub20].

Tables [MBLYR23]. Tail [YS24]. Tail-Optimal [YS24]. Tailed [SvKB^{+}20, YS24]. Tails [SvKB^{+}20]. TAO [PYL24]. Targeting [TMM20]. Tasks [GFS23, TKK21]. TCP [MSJ^{+}19, QWQ17]. Techniques [NKY^{+}18]. TeksDB [HKY^{+}19]. Telemetry [XJ22]. Telling [GYG^{+}18]. tenant [IYRR18]. Tensor [YX22]. term [SIN22]. Test [MMS^{+}23]. Testing [XLY^{+}22]. themed [WLZ^{+}22]. Theoretic [FZCL18]. Theory [KMM21, ZWT^{+}17]. Thermal [KKRK22]. Thermal-Aware [KKRK22]. Thickness [ABG^{+}19]. Thinking [PYL24]. Third [CEF^{+}20, KALB23, PBLC21]. Third-Party [CEF^{+}20, KALB23, PBLC21]. Thorough [CLE^{+}24]. Thread [JP23]. Threading [KB20]. Threats [ZYH^{+}19]. Threshold [HSV19]. Throughput [AAS23, GZB^{+}21, LC23, ZGB^{+}23]. Throw [CK23]. Time [AASS18, CCG21, HV21, HXW23, PSC19, PTD20, SvKB^{+}20, TL20, WYN18, KMM21, PHSB21, WMMC23]. Time-Sensitive [TL20]. Time-Series [PTD20]. Time-Varying [HXW23]. Times [GSHB19, WHBW22]. Timing [IYRR18]. Tokens [XWG^{+}21]. Tolerant [SDA18]. Tolerating [LGC^{+}18]. Tomographic [PST17]. Too [KB20]. Tool [RSJ24]. Toolkit [STGM21]. Topologies [MBvL18]. Topology [GZB^{+}21, WCX^{+}18, ZCW21]. Toxicity [ZRC^{+}22]. Trace [ALBS^{+}20]. Traceroute [MLHC20]. Traces [AGGS20]. Tracking [GWX^{+}20, LZGN22]. Trade [XWG^{+}21]. Tradeoff [ABG^{+}19]. Tradeoffs [TWFO20]. Traffic [ASV^{+}21, AGGS20, BSA^{+}19, BSA^{+}21, GSHB19, GSHBSW22, JLGJ^{+}24, LMS^{+}24, NPAP19, WCX^{+}18, XGS24, Yin17, ZWT^{+}17, ZZM^{+}19, ZLX^{+}23, ZWT^{+}18, ZTS18]. Traffic-driven [WCX^{+}18]. Training [CLE^{+}24, KKS22, LJG^{+}22]. Transformation [MYS^{+}22]. Transformer [CLE^{+}24]. Transient [AI19, SIS17]. Transit [PHSB21]. Transnational [ZMW^{+}20]. Transport [LMS^{+}24]. Trick [XWG^{+}21]. Troubleshooting [PZB^{+}23]. Truncation [RBT21]. Truth [HHP20]. TTL [KLCS19]. TTL-based [KLCS19]. Tuning [ZCM^{+}22]. Tuxedo [DAR^{+}21]. Two [CP18, WMX17]. Two-Sided [WMX17]. Type [CMM22].

Uncertain [ABG^{+}20, LYE19]. Uncertainty [HC18]. Understanding [CYG^{+}17, HWW^{+}20, JHWC21, LGL23, SVL20, SLWS20, WXW22, ZCW21, ZGN^{+}19]. Undirected [VSS22]. unfairness [ZMC24]. Unified [TSLG^{+}20]. Uniform [DMB^{+}22]. Unimodal [CPF20]. Uniswap [XWG^{+}21]. Unknown [FM22, YHW23]. Unleashing [BNS22]. Unreliable [PSS23, RCMW23]. Untrusted [LYQ^{+}22]. upon [GYSHB21]. URLs [CNK^{+}23]. Usefulness [SVL20]. User [CK23, KB20, PPL^{+}21]. User-Level [CK23, KB20]. Users [GFS23]. Using [BKO^{+}23, IYRR18, LJH^{+}23, PSC19, SZL19, WUNK17, YWB19, ZGB^{+}23, IBL^{+}19, JCS^{+}24, LPF^{+}22]. Utilities [FM22, WS17a]. Utility [TWFO20].

V5 [CCR21]. V6

[CCR22a, AGU22, CCR22b]. V7 [AGU23a, AGU23b, CFR23]. V8 [CFR24a, CFR24b]. Value [HKY^{+}19, SLJ19, SQ20]. Values [Gas17]. Variability [DDT^{+}22]. Variable [BBHB18]. Variation [LKS^{+}17, LGC^{+}18]. Varying [HXW23]. Vector [DE20, GFL^{+}22, MCS^{+}19]. Vehicle [SZL^{+}20, YWB19, ZHZ^{+}21]. Verification [LPJ23a]. via [ALBS^{+}20, AASS18, CKMP19, DE20, Gas17, HLH23, JJS17, LMS^{+}24, LZC^{+}23, NT19, ZLX^{+}23]. Video [BSA^{+}19, NGB^{+}21, WSC^{+}24]. View [CWW^{+}23, PJDQ20]. Violations [ZHZ^{+}21]. Virtual [LTS23]. Virtualization [WZT19, ZYH^{+}19]. Virtualized [AMA18, AIAR^{+}24, BKO^{+}23, NBB^{+}19, NKY^{+}18]. VirusTotal [CNK^{+}23]. Vision [RS22]. VM [PG17]. VMs [AI19]. VOD [LMS^{+}24]. Voltage [CYG^{+}17]. Volume [FM19]. Volume-based [FM19]. Voting [CL24]. VPNs [XVK22]. VR [TLL^{+}18]. vrfinder [MLHC20]. vs [CK23]. Vulnerability [QWQ17]. Vulnerable [ZAA^{+}24].

Waiting [HV21]. Walk [TS19]. Walks [DE20]. Wave [RS22]. Weak [YX22]. Weakly [WYN18]. Wear [LKJK20, ZGN^{+}19]. Weaving [HKY^{+}19]. Web [ASV^{+}21, LYL^{+}22, NB23, PSC19, ZRC^{+}22]. Webcams [SLWS20]. Wechat [ZTY^{+}21]. Weighted [IMMP20]. Welfare [FPCS22]. What-If [PSC19]. Where [BRC21]. Whittle [AdV18]. Who [AZG24, SVL20]. Wi [LC23, VPK18]. Wi-Fi [LC23, VPK18]. wide [AZG24]. Wiener [PSS23]. WiFi [ALBS^{+}20]. Wild [LGL23, SLWS20]. Wireless [AdV18, BZ18, KJM^{+}22]. WISEFUSE [MYS^{+}22]. Without [JIAM19, RVC^{+}19, BZ18, CPF20, PPL^{+}21]. WLANs [ANBK22]. Workflows [MYS^{+}22, RT24]. Working [NKY^{+}18]. Workload [GLH^{+}19, HLB^{+}23, HBV19, HNS19, MYS^{+}22]. Workload-DRAM [GLH^{+}19]. Workloads [AI19]. Works [AG22]. World [KALB23, OCW24]. Worst [YLIR24]. Write [ASME19, BKO^{+}23]. Write-Rationing [ASME19].

x [KJM^{+}22]. x-haul [KJM^{+}22]. Xaminer [RSJ24]. Xatu [NGB^{+}21].

Yes [DSR24]. YourAdvalue [PPL^{+}21].

Zero [WXHB21, WW20].

References

Anton:2020:IPH

[AAJV20] Elene Anton, Urtzi Ayesta, Matthieu Jonckheere, and Ina Maria Verloop. Improving the performance of heterogeneous data centers through redundancy. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(3):48:1–48:29, November 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3428333.

Addanki:2023:MNO

[AAS23] Vamsi Addanki, Chen Avin, and Stefan Schmid. Mars: Near-optimal throughput with shallow buffers in reconfigurable datacenter networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 7(1):2:1–2:??, March 2023. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3579312.

Agarwal:2018:MAT

[AASS18] Anish Agarwal, Muhammad Jehangir Amjad, Devavrat Shah, and Dennis Shen. Model agnostic time series analysis via matrix estimation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (3):40:1–40:39, December 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3287319.

Alijani:2019:STT

[ABG^{+}19] Reza Alijani, Siddhartha Banerjee, Sreenivas Gollapudi, Kostas Kollias, and Kamesh Munagala. The segmentationthickness tradeoff in online marketplaces. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):18:1–18:26, March 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3322205.3311089.

Alijani:2020:PMP

[ABG^{+}20] Reza Alijani, Siddhartha Banerjee, Sreenivas Gollapudi, Kamesh Munagala, and Kangning Wang. Predict and match: Prophet inequalities with uncertain supply. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):04:1– 04:23, May 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3379470.

Akbarzadeh:2024:HHB

[ADGF^{+}24] Negar Akbarzadeh, Sina Darabi, Atiyeh Gheibi-Fetrat, Amir Mirzaei, Mohammad Sadrosadati, and Hamid Sarbazi-Azad. H3DM: a high-bandwidth highcapacity hybrid 3D memory design for GPUs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(1):12:1–12:??, March 2024. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3639038.

Anand:2018:WIB

[AdV18] Arjun Anand and Gustavo de Veciana. A Whittle’s index based approach for QoE optimization in wireless networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):15:1–15:39, April 2018. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3179418.

Aral:2020:SCE

[AEKB20] Atakan Aral, Melike Erol-Kantarci, and Ivona Brandić. Staleness control for edge data analytics. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):38:1– 38:24, June 2020. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3392156.

Allmeier:2022:MFR

[AG22] Sebastian Allmeier and Nicolas Gast. Mean field and refined mean field approximations for heterogeneous systems: It works! Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 13:1–13:43, March 2022. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3508033.

Allmeier:2023:BRM

[AG23] Sebastian Allmeier and Nicolas Gast. Bias and refinement of multiscale mean field models. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):23:1– 23:??, March 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3579336.

Avin:2020:CTT

[AGGS20] Chen Avin, Manya Ghobadi, Chen Griner, and Stefan Schmid. On the complexity of traffic traces and implications. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):20:1– 20:29, May 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3379486.

Avrachenkov:2017:LCA

[AGS17] Konstantin Avrachenkov, Jasper Goseling, and Berksan Serbetci. A low-complexity approach to distributed cooperative caching with geographic constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):27:1– 27:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084465.

Avrachenkov:2022:PVN

[AGU22] Konstantin Avrachenkov, Phillipa Gill, and Bhuvan Urgaonkar. POMACS V6, N3 December 2022 editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(3):42:1–42:??, December 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3570602.

Avrachenkov:2023:PVNa

[AGU23a] Konstantin Avrachenkov, Phillipa Gill, and Bhuvan Urgaonkar. POMACS V7, N1, March 2023 editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1): 1:1–1:??, March 2023. CO-DEN ???? ISSN 2476-1249.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3579311.

Avrachenkov:2023:PVNb

[AGU23b] Konstantin Avrachenkov, Phillipa Gill, and Bhuvan Urgaonkar. POMACS V7, N2, June 2023 editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(2):28:1– 28:??, June 2023. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3589972.

Ambati:2019:OCE

[AI19] Pradeep Ambati and David Irwin. Optimizing the cost of executing mixed interactive and batch workloads on transient VMs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):28:1– 28:24, June 2019. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3341617.3326143.

Aslan:2024:FRA

[AIAR^{+}24] Fatih Aslan, George Iosifidis, Jose A. Ayala-Romero, Andres Garcia-Saavedra, and Xavier Costa-Perez. Fair resource allocation in virtualized O-RAN platforms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(1): 17:1–17:??, March 2024. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3639043.

Abanto-Leon:2020:SCL

[ALBS^{+}20] Luis Fernando Abanto-Leon, Andreas Bäuml, Gek Hong (Allyson) Sim, Matthias Hollick, and Arash Asadi. Stay connected, leave no trace: Enhancing security and privacy in WiFi via obfuscating radiometric fingerprints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(3):44:1–44:31, November 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3428329.

Aghajani:2017:PMA

[ALR17] Reza Aghajani, Xingjie Li, and Kavita Ramanan. The PDE method for the analysis of randomized load balancing networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):38:1–38:28, December 2017. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154497.

Abyaneh:2022:MMA

[ALZ22] Ali Hossein Abbasi Abyaneh, Maizi Liao, and Seyed Majid Zahedi. Malcolm: Multi-agent learning for cooperative load management at rack scale. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (3):59:1–59:??, December 2022. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3570611.

Ahmadian:2018:ECH

[AMA18] Saba Ahmadian, Onur Mutlu, and Hossein Asadi. ECI-Cache: a high-endurance and costefficient I/O caching scheme for virtualized platforms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (1):9:1–9:34, April 2018. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3179412.

Ayvasik:2023:PPE

[AMBK23] Serkut Ayvasik, Fidan Mehmeti, Edwin Babaians, and Wolfgang Kellerer. PEACH: Proactive and environment-aware channel state information prediction with depth images. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):24:1–24:??, March 2023. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3579450.

Amjad:2019:MMD

[AMSS19] Muhammad Amjad, Vishal Misra, Devavrat Shah, and Dennis Shen. mRSC: Multidimensional robust synthetic control. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):37:1– 37:27, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326152.

Asgari:2020:BSO

[AN20] Kamiar Asgari and Michael J.

Neely. Bregman-style online convex optimization with energy-harvesting constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(3):52:1–52:25, November 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3428337.

Aggarwal:2022:DLM

[ANBK22] Shivang Aggarwal, Srisai Karthik Neelamraju, Ajit Bhat, and Dimitrios Koutsonikolas. A detailed look at MIMO performance in 60 GHz WLANs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (2):38:1–38:??, June 2022. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3530904.

Ajdari:2022:EGO

[ARK^{+}22] Mohammadamin Ajdari, Patrick Raaf, Mostafa Kishani, Reza Salkhordeh, Hossein Asadi, and André Brinkmann. An enterprise-grade open-source data reduction architecture for all-flash storage systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (2):30:1–30:??, June 2022. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3530896.

Amjad:2017:CDE

[AS17] Muhammad J. Amjad and Devavrat Shah. Censored demand estimation in retail. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):31:1–31:28, December 2017. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154489.

Akram:2019:CGP

[ASME19] Shoaib Akram, Jennifer Sartor, Kathryn McKinley, and Lieven Eeckhout. Crystal gazer: Profile-driven writerationing garbage collection for hybrid memories. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1): 9:1–9:27, March 2019. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3322205.
3311080.

Akbari:2021:LBC

[ASV^{+}21] Iman Akbari, Mohammad A.

Salahuddin, Leni Ven, Noura Limam, Raouf Boutaba, Bertrand Mathieu, Stephanie Moteau, and Stephane Tuffin. A look behind the curtain: Traffic classification in an increasingly encrypted Web. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(1):04:1–04:26, February 2021. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3447382.

Ashok:2022:DDN

[ATN^{+}22] Sachin Ashok, Shubham Tiwari, Nagarajan Natarajan, Venkata N. Padmanabhan, and Sundararajan Sellamanickam. Data-driven network path simulation with iBox. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 6:1–6:26, March 2022. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3508026.

Anghel:2024:WGM

[AZG24] Radu Anghel, Yury Zhauniarovich, and Carlos Gañán. Who’s got my back? Measuring the adoption of an Internetwide BGP RTBH service. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8 (1):3:1–3:??, March 2024. CO-DEN ???? ISSN 2476-1249.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3639029.

AlMaruf:2023:MMD

[AZW^{+}23] Hasan Al Maruf, Yuhong Zhong, Hongyi Wang, Mosharaf Chowdhury, Asaf Cidon, and Carl Waldspurger. Memtrade: Marketplace for disaggregated memory clouds. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(2):41:1– 41:??, June 2023. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3589985.

Berger:2018:PBO

[BBHB18] Daniel S. Berger, Nathan Beckmann, and Mor Harchol-Balter. Practical bounds on optimal caching with variable object sizes. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2):32:1–32:38, June 2018. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3224427.

Bhattacharjee:2020:FLR

[BBS20] Rajarshi Bhattacharjee, Subhankar Banerjee, and Abhishek Sinha. Fundamental limits on the regret of online networkcaching. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):25:1– 25:31, June 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3392143.

Bonald:2017:PBF

[BCM17] Thomas Bonald, Céline Comte, and Fabien Mathieu. Performance of balanced fairness in resource pools: a recursive approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):41:1– 41:25, December 2017. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154500.

Bonald:2019:E

[BD19] Thomas Bonald and Nick Duffield. Editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (2):20:1–20:2, June 2019. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3341617.
3326134.

Berg:2017:TOP

[BDHB17] Benjamin Berg, Jan-Pieter Dorsman, and Mor Harchol-Balter. Towards optimality in parallel scheduling. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):40:1–40:30, December 2017. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154499.

Bakhshalipour:2023:AAS

[BG23]  Mohammad  Bakhshalipour

and Phillip B. Gibbons. Agents of autonomy: a systematic study of robotics on modern hardware. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(3):43:1–43:??, December 2023. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3626774.

Bitchebe:2023:GEB

[BKO^{+}23] Stella Bitchebe, Yves Kone, Pierre Olivier, Jalil Boukhobza, Yérom-David Bromberg, Daniel Hagimont, and Alain Tchana. GuaNary: Efficient buffer overflow detection in virtualized clouds using Intel EPT-based sub-page write protection support. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 7(3):56:1–56:??, December 2023. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3626787.

Bondorf:2017:QCD

[BNS17] Steffen Bondorf, Paul Nikolaus, and Jens B. Schmitt. Quality and cost of deterministic network calculus: Design and evaluation of an accurate and fast analysis. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):16:1– 16:??, June 2017. CODEN ????

ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084453.

Bouillard:2022:UPP

[BNS22] Anne Bouillard, Paul Nikolaus, and Jens Schmitt. Unleashing the power of paying multiplexing only once in stochastic network calculus. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2): 31:1–31:??, June 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3530897.

Bijlani:2021:WDM

[BRC21] Ashish Bijlani, Umakishore Ramachandran, and Roy Campbell. Where did my 256 GB go? A measurement analysis of storage consumption on smart mobile devices. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2): 28:1–28:28, June 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3460095.

Bronzino:2019:ISV

[BSA^{+}19] Francesco Bronzino, Paul Schmitt, Sara Ayoubi, Guilherme Martins, Renata Teixeira, and Nick Feamster. Inferring streaming video quality from encrypted traffic: Practical models and deployment experience. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):56:1–56:25, December 2019. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3366704.

Bronzino:2021:TRC

[BSA^{+}21] Francesco Bronzino, Paul Schmitt, Sara Ayoubi, Hyojoon Kim, Renata Teixeira, and Nick Feamster. Traffic refinery: Cost-aware data representation for machine learning on network traffic. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):40:1–40:24, December 2021. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3491052.

Borst:2018:DSM

[BZ18] Sem Borst and Martin Zubeldia. Delay scaling in manysources wireless networks without queue state information. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(2):34:1–34:45, June 2018. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3224429.

Bhuyan:2022:EEC

[BZY^{+}22] Sandeepa Bhuyan, Shulin Zhao, Ziyu Ying, Mahmut T. Kandemir, and Chita R. Das. End-to-end characterization of game streaming applications on mobile platforms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1):10:1–10:25, March 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508030.

Choi:2017:EDL

[CAJK17] Wonil Choi, Mohammad Arjomand, Myoungsoo Jung, and Mahmut Kandemir. Exploiting data longevity for enhancing the lifetime of flashbased storage class memory. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):21:1–21:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084458.

Casale:2017:API

[Cas17] Giuliano Casale. Accelerating performance inference over closed systems by asymptotic methods. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):8:1– 8:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084445.

Cardinaels:2024:MDS

[CBvL24] Ellen Cardinaels, Sem Borst, and Johan S. H. van Leeuwaarden. Multi-dimensional state space collapse in non-complete resource pooling scenarios.

Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 8(2):24:1–24:??, June 2024. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3656016.

Cuvelier:2021:SEP

[CCG21] Thibaut Cuvelier, Richard Combes, and Eric Gourdin. Statistically efficient, polynomial-time algorithms for combinatorial semi-bandits. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 5(1):09:1–09:31, February 2021. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3447387.

Carlsson:2021:PVN

[CCR21] Niklas Carlsson, Edith Cohen, and Philippe Robert. PO-MACS V5, N3, December 2021 editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):29:1, December 2021. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3491041.

Carlsson:2022:PVNa

[CCR22a] Niklas Carlsson, Edith Cohen, and Philippe Robert. PO-MACS V6, N1, March 2022 editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):1:1, March 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508021.

Carlsson:2022:PVNb

[CCR22b] Niklas Carlsson, Edith Cohen, and Philippe Robert. PO-MACS V6, N2, June 2022 editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2):24:1– 24:??, June 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3530890.

Cai:2020:TPD

[CEF^{+}20] Yang Cai, Federico Echenique, Hu Fu, Katrina Ligett, Adam Wierman, and Juba Ziani. Third-party data providers ruin simple mechanisms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (1):12:1–12:31, May 2020. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3379478.

Cayci:2019:LCR

[CES19] Semih Cayci, Atilla Eryilmaz, and R. Srikant. Learning to control renewal processes with bandit feedback. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):43:1– 43:32, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326158.

Chen:2023:GGA

[CF23] Canhui Chen and Zhixuan Fang. Gacha game analysis and design. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1): 6:1–6:??, March 2023. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3579438.

Ciucu:2023:PVN

[CFR23] Florin Ciucu, Giulia Fanti, and Rhonda Righter. POMACS V7, N3, December 2023 editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(3):42:1– 42:??, December 2023. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3626773.

Ciucu:2024:PVNa

[CFR24a] Florin Ciucu, Giulia Fanti, and Rhonda Righter. POMACS V8, N1, March 2024 editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8 (1):1:1–1:??, March 2024. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3639027.

Ciucu:2024:PVNb

[CFR24b] Florin Ciucu, Giulia Fanti, and Rhonda Righter. POMACS V8, N2, June 2024 editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8 (2):18:1–18:??, June 2024. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3656013.

Chaintreau:2017:E

[CGZ17] Augustin Chaintreau, Leana Golubchik, and Zhi-Li Zhang. Editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):1:1– 1:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3105875.

Cui:2022:SAC

[CHT^{+}22] Shuang Cui, Kai Han, Jing Tang, He Huang, Xueying Li, and Zhiyu Li. Streaming algorithms for constrained submodular maximization. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (3):54:1–54:??, December 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3570615.

Chang:2019:LBN

[CJC^{+}19] Yiyang Chang, Chuan Jiang, Ashish Chandra, Sanjay Rao, and Mohit Tawarmalani. Lancet: Better network resilience by designing for pruned failure sets. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 3(3):49:1–49:26, December 2019. CODEN ????

ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3366697.

Cai:2023:KVU

[CK23] Peter Cai and Martin Karsten. Kernel vs. user-level networking: Don’t throw out the stack with the interrupts. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7 (3):49:1–49:??, December 2023. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3626780.

Cullina:2019:PRE

[CKMP19] Daniel Cullina, Negar Kiyavash, Prateek Mittal, and H. Vincent Poor. Partial recovery of Erdős–Rényi graph alignment via k-core alignment. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):54:1–54:21, December 2019. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3366702.

Chu:2024:FLQ

[CL24] Tianyue Chu and Nikolaos Laoutaris. FedQV: Leveraging quadratic voting in federated learning. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(2):22:1– 22:??, June 2024. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3656006.

Cheng:2024:TCA

[CLE^{+}24] Scott Cheng, Jun-Liang Lin, Murali Emani, Siddhisanket Raskar, Sam Foreman, Zhen Xie, Venkatram Vishwanath, and Mahmut Taylan Kandemir. Thorough characterization and analysis of large transformer model training atscale. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 8(1):8:1–8:??, March 2024. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3639034.

Chen:2021:SDC

[CLS^{+}21] Weimin Chen, Xinran Li, Yuting Sui, Ningyu He, Haoyu Wang, Lei Wu, and Xiapu Luo. SADPonzi: Detecting and characterizing Ponzi schemes in Ethereum smart contracts. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2):26:1– 26:30, June 2021. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3460093.

Chen:2017:FGE

[CMH^{+}17] Xiaomeng Chen, Jiayi Meng, Y. Charlie Hu, Maruti Gupta, Ralph Hasholzner, Venkatesan Nallampatti Ekambaram, Ashish Singh, and Srikathyayani Srikanteswara. A fine-grained event-based modem power model for enabling in-depth modem energy drain analysis. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 1(2):45:1–45:28, December 2017. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3154504.

Chen:2022:SBC

[CMM22] Zaiwei Chen, Shancong Mou, and Siva Theja Maguluri. Stationary behavior of constant stepsize SGD type algorithms: an asymptotic characterization. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1):19:1–19:24, March 2022. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3508039.

Choo:2023:LSS

[CNK^{+}23] Euijin Choo, Mohamed Nabeel, Doowon Kim, Ravindu De Silva, Ting Yu, and Issa Khalil. A large scale study and classification of VirusTotal reports on phishing and malware URLs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 7(3):59:1–59:??, December 2023. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3626790.

Cao:2017:DEC

[CNW^{+}17] Yi Cao, Javad Nejati, Muhammad Wajahat, Aruna Balasubramanian, and Anshul Gandhi. Deconstructing the energy consumption of the mobile page load. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):6:1– 6:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084443.

Ciucu:2018:TEK

[CP18] Florin Ciucu and Felix Poloczek.

Two extensions of Kingman’s GI/G/1 bound. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(3):43:1–43:33, December 2018. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3287322.

Combes:2020:UBC

[CPF20] Richard Combes, Alexandre Proutière, and Alexandre Fauquette. Unimodal bandits with continuous arms: Orderoptimal regret without smoothness. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):14:1–14:28, May 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3379480.

Ciucu:2019:QLD

[CPR19] Florin Ciucu, Felix Poloczek, and Amr Rizk. Queue and loss distributions in finitebuffer queues. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):31:1– 31:29, June 2019. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3341617.3326146.

Chen:2017:DSM

[CSX17] Yudong Chen, Lili Su, and Jiaming Xu. Distributed statistical machine learning in adversarial settings: Byzantine gradient descent. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):44:1–44:25, December 2017. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3154503.

Combes:2019:CEE

[CT19] Richard Combes and Mikael Touati. Computationally efficient estimation of the spectral gap of a Markov chain. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (1):7:1–7:21, March 2019. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3322205.
                                        3311078.

Chiesa:2023:NMM

[CV23] Marco Chiesa and Fábio L. Verdi. Network monitoring on multi-pipe switches. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):

8:1–8:??, March 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3579321.

Carisimo:2023:HEV

[CWW^{+}23] Esteban Carisimo, Caleb J.

Wang, Mia Weaver, Fabián E. Bustamante, and Paul Barford. A hop away from everywhere: a view of the intercontinental long-haul infrastructure. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7 (3):47:1–47:??, December 2023. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3626778.

Comden:2019:OOC

[CYC^{+}19] Joshua Comden, Sijie Yao, Niangjun Chen, Haipeng Xing, and Zhenhua Liu. Online optimization in cloud resource provisioning: Predictions, regrets, and algorithms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):16:1–16:30, March 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3322205.3311087.

Chang:2017:URV

[CYG^{+}17] Kevin K. Chang, A. Giray Yaălikçi, Saugata Ghose, Aditya Agrawal, Niladrish Chatterjee, Abhijith Kashyap, Donghyuk Lee, Mike O’Connor, Hasan Hassan, and Onur Mutlu. Understanding reducedvoltage operation in modern DRAM devices: Experimental characterization, analysis, and mechanisms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):10:1– 10:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

                                        /dl.acm.org/citation.cfm?
                                        id=3084447.

Das:2021:TMS

[DAR^{+}21] Sourav Das, Nitin Awathare, Ling Ren, Vinay J. Ribeiro, and Umesh Bellur. Tuxedo: Maximizing smart contract computation in PoW blockchains. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 5(3):41:1–41:30, December 2021. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3491053.

Doan:2017:CRD

[DBS17] Thinh T. Doan, Carolyn L. Beck, and R. Srikant. On the convergence rate of distributed gradient methods for finite-sum optimization under communication delays. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):37:1–37:27, December 2017. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3154496.

Dai:2019:ACL

[DCKG19] Osman Emre Dai, Daniel Cullina, Negar Kiyavash, and Matthias Grossglauser. Analysis of a canonical labeling algorithm for the alignment of correlated Erdős–Rényi graphs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):36:1– 36:25, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326151.

DeSensi:2022:NCI

[DDT^{+}22] Daniele De Sensi, Tiziano De Matteis, Konstantin Taranov, Salvatore Di Girolamo, Tobias Rahn, and Torsten Hoefler. Noise in the clouds: Influence of network performance variability on application scalability. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (3):49:1–49:??, December 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3570609.

Doshi:2020:FVA

[DE20] Vishwaraj Doshi and Do Young Eun. Fiedler vector approximation via interacting random walks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):01:1– 01:28, May 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3379502.

Darabi:2022:NFS

[DMB^{+}22] Sina Darabi, Negin Mahani, Hazhir Baxishi, Ehsan Yousefzadeh-Asl-Miandoab, Mohammad Sadrosadati, and Hamid Sarbazi-Azad. NURA: a framework for supporting non-uniform resource accesses in GPUs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1):16:1–16:27, March 2022. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3508036.

Dozier:2024:AFN

[DSR24] Kahlil Dozier, Loqman Salamatian, and Dan Rubenstein. Analysis of false negative rates for recycling Bloom filters (yes, they happen!). Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(2):21:1– 21:??, June 2024. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3656005.

Duran:2018:AOC

[DV18] Santiago Duran and Ina Maria Verloop. Asymptotic optimal control of Markov-modulated restless bandits. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (1):7:1–7:25, April 2018. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3179410.

Eliav:2018:BGD

[EC18] Buchnik Eliav and Edith Cohen. Bootstrapped graph diffusions: Exposing the power of nonlinearity. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):10:1–10:19, April 2018. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3179413.

Foerster:2021:IDD

[FKP^{+}21] Klaus-Tycho Foerster, Janne H.

Korhonen, Ami Paz, Joel Rybicki, and Stefan Schmid. Input-dynamic distributed algorithms for communication networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(1):06:1–06:33, February 2021. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3447384.

Fu:2019:FLV

[FM19] Xinzhe Fu and Eytan Modiano. Fundamental limits of volume-based network DoS attacks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):50:1– 50:36, December 2019. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366698.

Fu:2022:JLC

[FM22] Xinzhe Fu and Eytan Modiano. Joint learning and control in stochastic queueing networks with unknown utilities. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (3):58:1–58:??, December 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3570619.

Faw:2022:LMW

[FPCS22] Matthew Faw, Orestis Papadigenopoulos, Constantine Caramanis, and Sanjay Shakkottai. Learning to maximize welfare with a reusable resource. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2):27:1– 27:??, June 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3530893.

Fleder:2019:FAD

[FS19] Michael Fleder and Devavrat Shah. Forecasting with alternative data. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):46:1– 46:29, December 2019. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366694.

Fleder:2020:KWY

[FS20] Michael Fleder and Devavrat Shah. I know what you bought at Chipotle for $9.81 by solving a linear inverse problem. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):47:1–47:17, November 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3428332.

Fanti:2018:DLC

[FVB^{+}18] Giulia Fanti, Shaileshh Bojja Venkatakrishnan, Surya Bakshi, Bradley Denby, Shruti Bhargava, Andrew Miller, and Pramod Viswanath. Dandelion++: Lightweight cryptocurrency networking with formal anonymity guarantees. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(2):29:1–29:35, June 2018. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3224424.

Formby:2017:CSP

[FWB17] David Formby, Anwar Walid, and Raheem Beyah. A case study in power substation network dynamics. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):19:1– 19:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084456.

Freeman:2018:DPS

[FZCL18] Rupert Freeman, Seyed Majid Zahedi, Vincent Conitzer, and Benjamin C. Lee. Dynamic proportional sharing: a game-theoretic approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (1):3:1–3:36, April 2018. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3179406.

Gozlan:2023:GCB

[GAES23] Itamar Gozlan, Chen Avin, Gil Einziger, and Gabriel Scalosub. Go-to-controller is better: Efficient and optimal LPM caching with splicing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1): 10:1–10:??, March 2023. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3579441.

Gast:2017:EVE

[Gas17] Nicolas Gast. Expected values estimated via mean-field approximation are 1/N-accurate. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 1(1):17:1–17:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

                                        /dl.acm.org/citation.cfm?
                                        id=3084454.

Goldsztajn:2024:SSS

[GBvL24] Diego Goldsztajn, Sem C. Borst, and Johan S. H. van Leeuwaarden. Server saturation in skewed networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8 (2):26:1–26:??, June 2024. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3656018.

Giannoula:2022:STE

[GFL^{+}22] Christina Giannoula, Ivan Fernandez, Juan Gómez Luna, Nectarios Koziris, Georgios Goumas, and Onur Mutlu. SparseP: Towards efficient sparse matrix vector multiplication on real processing-in-memory architectures. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1):21:1–21:49, March 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508041.

Goldberg:2023:BTU

[GFS23] Alexander Goldberg, Giulia Fanti, and Nihar B. Shah. Batching of tasks by users of pseudonymous forums: Anonymity compromise and protection. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):22:1–22:??, March 2023. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3579335.

Gibbens:2017:HND

[GGYZ17] Mathias Gibbens, Chris Gniady, Lei Ye, and Beichuan Zhang. Hadoop on named data networking: Experience and results. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):2:1– 2:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

                                        /dl.acm.org/citation.cfm?
                                        id=3084439.

Giannoula:2023:DAS

[GHT^{+}23] Christina Giannoula, Kailong Huang, Jonathan Tang, Nectarios Koziris, Georgios Goumas, Zeshan Chishti, and Nandita Vijaykumar. DaeMon: Architectural support for efficient data movement in fully disaggregated systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):16:1–16:??, March 2023. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3579445.

Ghahani:2020:DCH

[GKK20] Seyed Armin Vakil Ghahani, Mahmut Taylan Kandemir, and Jagadish B. Kotra. DSM: a case for hardware-assisted merging of DRAM rows with same content. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):33:1– 33:26, June 2020. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3392151.

Gandhi:2021:E

[GKW21] Anshul Gandhi, Negar Kiyavash, and Jia Wang. Editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2):13:1, June 2021. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3466793.

Ghose:2019:DCW

[GLH^{+}19] Saugata Ghose, Tianshi Li, Nastaran Hajinazar, Damla Senol Cali, and Onur Mutlu. Demystifying complex workload-DRAM interactions: an experimental study. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):60:1–60:50, December 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3366708.

Grosof:2019:LBG

[GSHB19] Isaac Grosof, Ziv Scully, and Mor Harchol-Balter. Load balancing guardrails: Keeping your heavy traffic on the road to low response times. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2): 42:1–42:31, June 2019. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3341617.
3326157.

Grosof:2022:OSM

[GSHBSW22] Isaac Grosof, Ziv Scully, Mor Harchol-Balter, and Alan Scheller-Wolf. Optimal scheduling in the multiserver-job model under heavy traffic. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (3):51:1–51:??, December 2022. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3570612.

Gopalan:2020:SSB

[GSWV20] Aditya Gopalan, Abishek Sankararaman, Anwar Walid, and Sriram Vishwanath. Stability and scalability of blockchain systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):35:1– 35:35, June 2020. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3392153.

Gong:2017:QPS

[GTL^{+}17] Long Gong, Paul Tune, Liang Liu, Sen Yang, and Jun (Jim) Xu. Queue-proportional sampling: A better approach to crossbar scheduling for inputqueued switches. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):3:1– 3:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

                                        /dl.acm.org/citation.cfm?
                                        id=3084440.

Gast:2017:RMF

[GV17] Nicolas Gast and Benny Van Houdt. A refined mean field approximation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):33:1– 33:28, December 2017. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154491.

Gast:2024:ASI

[GV24] Nicolas Gast and Benny Van Houdt. Approximations to study the impact of the service discipline in systems with redundancy. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(1):14:1– 14:??, March 2024. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3639040.

Gao:2020:TCC

[GWX^{+}20] Bingyu Gao, Haoyu Wang, Pengcheng Xia, Siwei Wu, Yajin Zhou, Xiapu Luo, and Gareth Tyson. Tracking counterfeit cryptocurrency end-to-end. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(3):50:1–50:28, November 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3428335.

Ghose:2018:WYD

[GYG^{+}18] Saugata Ghose, Abdullah Giray Yaglikçi, Raghav Gupta, Donghyuk Lee, Kais Kudrolli, William X. Liu, Hasan Hassan, Kevin K. Chang, Niladrish Chatterjee, Aditya Agrawal, Mike O’Connor, and Onur Mutlu. What your DRAM power models are not telling you: Lessons from a detailed experimental study. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (3):38:1–38:41, December 2018. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3224419.

Grosof:2021:NSI

[GYSHB21] Isaac Grosof, Kunhe Yang, Ziv Scully, and Mor Harchol-Balter. Nudge: Stochastically improving upon FCFS. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (2):21:1–21:29, June 2021. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3460088.

Griner:2021:CPC

[GZB^{+}21] Chen Griner, Johannes Zerwas, Andreas Blenk, Manya Ghobadi, Stefan Schmid, and Chen Avin. Cerberus: The power of choices in datacenter topology design — a throughput perspective. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):38:1–38:33, December 2021. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3491050.

Hellemans:2019:PAW

[HBV19] Tim Hellemans, Tejas Bodas, and Benny Van Houdt. Performance analysis of workload dependent load balancing policies. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):35:1– 35:35, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326150.

Hoffmann:2018:CUC

[HC18] Jessica Hoffmann and Constantine Caramanis. The cost of uncertainty in curing epidemics. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2):31:1– 31:33, June 2018. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3224426.

Hoffmann:2019:LGN

[HC19] Jessica Hoffmann and Constantine Caramanis. Learning graphs from noisy epidemic cascades. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):40:1– 40:34, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326155.

Han:2021:AAS

[HCZ^{+}21] Kai Han, Shuang Cui, Tianshuai Zhu, Enpei Zhang, Benwei Wu, Zhizhuo Yin, Tong Xu, Shaojie Tang, and He Huang. Approximation algorithms for submodular data summarization with a knapsack constraint. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(1):05:1– 05:31, February 2021. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3447383.

Huang:2023:MMM

[HHM^{+}23] Jintao Huang, Ningyu He, Kai Ma, Jiang Xiao, and Haoyu Wang. Miracle or mirage? A measurement study of NFT rug pulls. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(3):51:1– 51:??, December 2023. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3626782.

Hazimeh:2020:MGT

[HHP20] Ahmad Hazimeh, Adrian Herrera, and Mathias Payer. Magma: a ground-truth fuzzing benchmark. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):49:1–49:29, November 2020. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3428334.

Han:2019:TWD

[HKY^{+}19] Youil Han, Bryan S. Kim, Jeseong Yeon, Sungjin Lee, and Eunji Lee. TeksDB: Weaving data structures for a high-performance key–value store. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 3(1):8:1–8:23, March 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3322205.3311079.

Hanafy:2023:CLC

[HLB^{+}23] Walid A. Hanafy, Qianlin Liang, Noman Bashir, David Irwin, and Prashant Shenoy. CarbonScaler: Leveraging cloud workload elasticity for optimizing carbon-efficiency. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 7(3):57:1–57:??, December 2023. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3626788.

Huang:2023:OCP

[HLH23] Yudi Huang, Yilei Lin, and Ting He. Optimized crosspath attacks via adversarial reconnaissance. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(3):58:1– 58:??, December 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3626789.

Hsu:2023:FLI

[HLP23] Amanda Hsu, Frank Li, and Paul Pearce. Fiat Lux: Illuminating IPv6 apportionment with different datasets. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):21:1–21:??, March 2023. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3579334.

Henzinger:2019:EDW

[HNS19] Monika Henzinger, Stefan Neumann, and Stefan Schmid. Efficient distributed workload (re- )Embedding. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):13:1–13:38, March 2019. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3322205.3311084.

Horvath:2019:MFA

[HSV19] Illés Antal Horváth, Ziv Scully, and Benny Van Houdt. Mean field analysis of join-below-threshold load balancing for resource sharing servers. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):57:1–57:21, December 2019. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3366705.

Hellemans:2018:PDC

[HV18] Tim Hellemans and Benny Van Houdt. On the power-of-d-choices with least loaded server selection. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2):27:1– 27:22, June 2018. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3224422.

Hellemans:2021:MWT

[HV21] Tim Hellemans and Benny Van Houdt. Mean waiting time in large-scale and critically loaded power of d load balancing systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2): 19:1–19:34, June 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3460086.

Huang:2020:UMB

[HWW^{+}20] Yuheng Huang, Haoyu Wang, Lei Wu, Gareth Tyson, Xiapu Luo, Run Zhang, Xuanzhe Liu, Gang Huang, and Xuxian Jiang. Understanding (mis)Behavior on the EO-SIO blockchain. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):37:1– 37:28, June 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3392155.

Hong:2023:NOS

[HXW23] Yige Hong, Qiaomin Xie, and Weina Wang. Near-optimal stochastic bin-packing in large service systems with timevarying item sizes. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7 (3):48:1–48:??, December 2023. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3626779.

Ivkin:2019:KWY

[IBL^{+}19] Nikita Ivkin, Ran Ben Basat, Zaoxing Liu, Gil Einziger, Roy Friedman, and Vladimir Braverman. I know what you did last summer: Network monitoring using interval queries. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):61:1–61:28, December 2019. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3366709.

Iqbal:2021:DCG

[IKS21] Hassan Iqbal, Ayesha Khalid, and Muhammad Shahzad. Dissecting cloud gaming performance with DECAF. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):31:1–31:27, December 2021. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3491043.

Iyer:2023:ABA

[ILL^{+}23] Venkatraman Iyer, Sungho Lee, Semun Lee, Juitem Joonwoo Kim, Hyunjun Kim, and Youngjae Shin. Automated backend allocation for multimodel, on-device AI inference. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 7(3):62:1–62:??, December 2023. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3626793.

Im:2020:DWF

[IMMP20] Sungjin Im, Benjamin Moseley, Kamesh Munagala, and Kirk Pruhs. Dynamic weighted fairness with minimal disruptions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):19:1–19:18, May 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3379485.

Izhikevich:2024:DLS

[ITI^{+}24] Liz Izhikevich, Manda Tran, Katherine Izhikevich, Gautam Akiwate, and Zakir Durumeric. Democratizing LEO satellite network measurement. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(1):13:1–13:??, March 2024. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3639039.

Islam:2018:WSL

[IYRR18] Mohammad A. Islam, Luting Yang, Kiran Ranganath, and Shaolei Ren. Why some like it loud: Timing power attacks in multi-tenant data centers using an acoustic side channel. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (1):6:1–6:33, April 2018. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3179409.

Jonatan:2024:SLP

[JCS^{+}24] Gilbert Jonatan, Haeyoon Cho, Hyojun Son, Xiangyu Wu, Neal Livesay, Evelio Mora, Kaustubh Shivdikar, José L. Abellán, Ajay Joshi, David Kaeli, and John Kim. Scalability limitations of processingin-memory using real system evaluations. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(1): 5:1–5:??, March 2024. CO-DEN ???? ISSN 2476-1249.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3639046.

Jin:2021:UPG

[JHWC21] Lin Jin, Shuai Hao, Haining Wang, and Chase Cotton. Understanding the practices of global censorship through accurate, end-to-end measurements. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):43:1– 43:25, December 2021. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3491055.

Jose:2019:DAC

[JIAM19] Lavanya Jose, Stephen Ibanez, Mohammad Alizadeh, and Nick McKeown. A distributed algorithm to calculate maxmin fair rates without perflow state. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):21:1– 21:42, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326135.

Ju:2017:HLS

[JJS17] Xiaoen Ju, Hani Jamjoom, and Kang G. Shin. Hieroglyph: Locally-sufficient graph processing via compute–sync– merge. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):9:1– 9:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084446.

Jain:2018:QEC

[JKR18] Akshay Jain, Mahmoud Khairy, and Timothy G. Rogers. A quantitative evaluation of contemporary GPU simulation methodology. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2):35:1– 35:28, June 2018. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3224430.

Jiang:2024:NND

[JLGJ^{+}24] Xi Jiang, Shinan Liu, Aaron Gember-Jacobson, Arjun Nitin Bhagoji, Paul Schmitt, Francesco Bronzino, and Nick Feamster. NetDiffusion: Network data augmentation through protocol-constrained traffic generation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(1):11:1–11:??, March 2024. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3639037.

Jiang:2018:CSM

[JNTG18] Bo Jiang, Philippe Nain, Don Towsley, and Saikat Guha. On a class of stochastic multilayer networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1): 18:1–18:25, April 2018. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3179421.

Jin:2023:SSL

[JP23] Wangkai Jin and Xiangjun Peng. SLITS: Sparsitylightened intelligent thread scheduling. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):3:1– 3:??, March 2023. CO-DEN ???? ISSN 2476-1249.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3579436.

Kumar:2023:EOP

[KALB23] Rashna Kumar, Sana Asif, Elise Lee, and Fabian E. Bustamante. Each at its own pace: Third-party dependency and centralization around the world. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 7(1):4:1–4:??, March 2023. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3579437.

Karsten:2020:ULT

[KB20] Martin Karsten and Saman Barghi. User-level threading: Have your cake and eat it too. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):17:1– 17:30, May 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3379483.

Kuo:2020:SCH

[KCMX20] Hsuan-Chi Kuo, Jianyan Chen, Sibin Mohan, and Tianyin Xu. Set the configuration for the heart of the OS: On the practicality of operating system kernel debloating. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1): 03:1–03:27, May 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3379469.

Kuhnle:2018:NRL

[KCT18] Alan Kuhnle, Victoria G.

Crawford, and My T. Thai. Network resilience and the length-bounded multicut problem: Reaching the dynamic billion-scale with guarantees. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(1):4:1–4:26, April 2018. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3179407.

Kadota:2022:SRP

[KJM^{+}22] Igor Kadota, Dror Jacoby, Hagit Messer, Gil Zussman, and Jonatan Ostrometzky. Switching in the rain: Predictive wireless x-haul network reconfiguration. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(3):55:1–55:??, December 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3570616.

Kao:2022:FDA

[KKP^{+}22] Sheng-Chun Kao, Hyoukjun Kwon, Michael Pellauer, Angshuman Parashar, and Tushar Krishna. A formalism of DNN accelerator flexibility. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2): 41:1–41:??, June 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3530907.

Khadirsharbiyani:2022:DCG

[KKRK22] Soheil Khadirsharbiyani, Jagadish Kotra, Karthik Rao, and Mahmut Kandemir. Data convection: a GPU-driven case study for thermal-aware data placement in 3D DRAMs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (1):7:1–7:25, March 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508027.

K:2022:CPA

[KKS22] Prashanthi S. K, Sai Anuroop Kesanapalli, and Yogesh Simmhan. Characterizing the performance of accelerated Jetson edge devices for training deep learning models. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(3):44:1–44:??, December 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3570604.

Karakoy:2019:AAA

[KKT^{+}19] Mustafa Karakoy, Orhan Kislal, Xulong Tang, Mahmut Taylan Kandemir, and Meenakshi Arunachalam. Architectureaware approximate computing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):38:1–38:24, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326153.

Kumar:2019:TBA

[KLCS19] Dhruv Kumar, Jian Li, Abhishek Chandra, and Ramesh Sitaraman. A TTL-based approach for data aggregation in geo-distributed streaming analytics. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):29:1– 29:27, June 2019. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3341617.3326144.

Kinnear:2021:RTB

[KMM21] Ryan J. Kinnear, Ravi R. Mazumdar, and Peter Marbach. Real-time bidding for time constrained impression contracts in first and second price auctions — theory and algorithms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):37:1–37:37, December 2021. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3491049.

Kumar:2023:SCD

[KSZ23] Adithya Kumar, Anand Sivasubramaniam, and Timothy Zhu. SplitRPC: a control + data path splitting RPC stack for ML inference serving. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7 (2):30:1–30:??, June 2023. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3589974.

Kielanski:2021:AIS

[KV21] Grzegorz Kielanski and Benny Van Houdt. On the asymptotic insensitivity of the supermarket model in processor sharing systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2): 22:1–22:28, June 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3460089.

Liu:2023:FLW

[LC23] Ruofeng Liu and Nakjung Choi. A first look at Wi-Fi 6 in action: Throughput, latency, energy efficiency, and security. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):25:1–25:??, March 2023. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3579451.

Li:2021:IAC

[LCS^{+}21] Tongxin Li, Yue Chen, Bo Sun, Adam Wierman, and Steven H. Low. Information aggregation for constrained online control. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2):18:1–18:35, June 2021. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3460085.

Lechowicz:2023:OPR

[LCZ^{+}23] Adam Lechowicz, Nicolas Christianson, Jinhang Zuo, Noman Bashir, Mohammad Hajiesmaili, Adam Wierman, and Prashant Shenoy. The online pause and resume problem: Optimal algorithms and an application to carbon-aware load shifting. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(3):45:1–45:??, December 2023. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3626776.

Liu:2024:LOC

[LF24] Qingsong Liu and Zhixuan Fang. Learning the optimal control for evolving systems with converging dynamics. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(2):23:1–23:??, June 2024. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3656007.

Luo:2018:INF

[LGC^{+}18] Yixin Luo, Saugata Ghose, Yu Cai, Erich F. Haratsch, and Onur Mutlu. Improving 3D NAND flash memory lifetime by tolerating early retention loss and process variation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (3):37:1–37:48, December 2018. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3224432.

Li:2017:SYE

[LGH^{+}17] Lingda Li, Robel Geda, Ari B. Hayes, Yanhao Chen, Pranav Chaudhari, Eddy Z. Zhang, and Mario Szegedy. A simple yet effective balanced edge partition model for parallel computing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):14:1– 14:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084451.

Luo:2022:DRM

[LGK22] Yuwei Luo, Varun Gupta, and Mladen Kolar. Dynamic regret minimization for control of non-stationary linear dynamical systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 9:1–9:72, March 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508029.

Li:2023:TUC

[LGL23] Kai Li, Shixuan Guan, and Darren Lee. Towards understanding and characterizing the arbitrage bot scam in the wild. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(3):52:1– 52:??, December 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3626783.

Lin:2020:OOP

[LGW20] Yiheng Lin, Gautam Goel, and Adam Wierman. Online optimization with predictions and non-convex losses. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1): 18:1–18:32, May 2020. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3379484.

Liu:2022:FPS

[LJG^{+}22] Yunzhuo Liu, Bo Jiang, Tian Guo, Zimeng Huang, Wenhao Ma, Xinbing Wang, and Chenghu Zhou. FuncPipe: a pipelined serverless framework for fast and cost-efficient training of deep learning models. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(3):47:1–47:??, December 2022. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3570607.

Lin:2023:TAD

[LJH^{+}23] Jiaxin Lin, Tao Ji, Xiangpeng Hao, Hokeun Cha, Yanfang Le, Xiangyao Yu, and Aditya Akella. Towards accelerating data intensive Application’s shuffle process using SmartNICs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(2):36:1– 36:??, June 2023. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3589980.

Lee:2019:NMM

[LKE19] Chul-Ho Lee, Min Kang, and Do Young Eun. Non-Markovian Monte Carlo on directed graphs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):15:1–15:31, March 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3322205.3311086.

Liu:2020:CNA

[LKJK20] Chun-Yi Liu, Jagadish Kotra, Myoungsoo Jung, and Mahmut Taylan Kandemir. Centaur: a novel architecture for reliable, low-wear, high-density 3D NAND storage. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2): 28:1–28:25, June 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392146.

Lee:2017:DIL

[LKS^{+}17] Donghyuk Lee, Samira Khan, Lavanya Subramanian, Saugata Ghose, Rachata Ausavarungnirun, Gennady Pekhimenko, Vivek Seshadri, and Onur Mutlu. Design-induced latency variation in modern DRAM chips: Characterization, analysis, and latency reduction mechanisms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):26:1– 26:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084464.

Liang:2018:MQL

[LM18] Qingkai Liang and Eytan Modiano. Minimizing queue length regret under adversarial network models. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):11:1–11:32, April 2018. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3179414.

Lin:2020:FDA

[LML^{+}20] Fred Lin, Keyur Muzumdar, Nikolay Pavlovich Laptev, Mihai-Valentin Curelea, Seunghak Lee, and Sriram Sankar. Fast dimensional analysis for root cause investigation in a large-scale service environment. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):31:1–31:23, June 2020. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3392149.

Lo:2024:SVT

[LMS^{+}24] Chi-Jen (Roger) Lo, Mahesh K. Marina, Nishanth Sastry, Kai Xu, Saeed Fadaei, and Yong Li. Shrinking VOD traffic via Rényi–entropic optimal transport. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 8(1):7:1–7:??, March 2024. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3639033.

Lin:2022:COO

[LMSC22] Qiulin Lin, Yanfang Mo, Junyan Su, and Minghua Chen. Competitive online optimization with multiple inventories: a divide-and-conquer approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2):36:1– 36:??, June 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3530902.

Lange:2021:OOA

[LNY21] Tomer Lange, Joseph (Seffi)

Naor, and Gala Yadgar. Offline and online algorithms for SSD management. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):33:1–33:28, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491045.

Li:2022:SAH

[LPF^{+}22] Lingda Li, Santosh Pandey, Thomas Flynn, Hang Liu, Noel Wheeler, and Adolfy Hoisie. SimNet: Accurate and highperformance computer architecture simulation using deep learning. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2):25:1– 25:??, June 2022. CODEN ????

ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3530891.

Lindstaahl:2023:CPD

[LPJ23a] Simon Lindståhl, Alexandre Proutiere, and Andreas Johnsson. Change point detection with adaptive measurement schedules for network performance verification. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7 (3):53:1–53:??, December 2023. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3626784.

Liu:2023:AAP

[LPJ23b] Hongyuan Liu, Sreepathi Pai, and Adwait Jog. Asynchronous automata processing on GPUs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):27:1–27:??, March 2023. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3579453.

Li:2021:OCN

[LSNI21] Yuanyuan Li, Tareq Si Salem, Giovanni Neglia, and Stratis Ioannidis. Online caching networks with adversarial guarantees. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):35:1– 35:39, December 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491047.

Liu:2023:SFF

[LSX^{+}23] Yi Liu, Shouqian Shi, Minghao Xie, Heiner Litz, and Chen Qian. Smash: Flexible, fast, and resource-efficient placement and lookup of distributed storage. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(2): 33:1–33:??, June 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3589977.

Liu:2022:DPP

[LT22] Mozhengfu Liu and Xueyan Tang. Dynamic bin packing with predictions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (3):45:1–45:??, December 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3570605.

Lyu:2023:MMM

[LTS23] Minzhao Lyu, Rahul Dev Tripathi, and Vijay Sivaraman. MetaVRadar: Measuring metaverse virtual reality network activity. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(3):55:1–55:??, December 2023. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3626786.

London:2019:LCD

[LVW19] Palma London, Shai Vardi, and Adam Wierman. Logarithmic communication for distributed optimization in multiagent systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):48:1– 48:29, December 2019. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3366696.

Liu:2019:PCL

[LYE19] Ran Liu, Edmund Yeh, and Atilla Eryilmaz. Proactive caching for low access-delay services under uncertain predictions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):2:1– 2:46, March 2019. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3322205.
                                        3311073.

Lu:2021:OPD

[LYJ^{+}21] Bingqian Lu, Jianyi Yang, Weiwen Jiang, Yiyu Shi, and Shaolei Ren. One proxy device is enough for hardware-aware neural architecture search. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):34:1–34:34, December 2021. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3491046.

Liu:2022:FSI

[LYL^{+}22] Wei Liu, Xinlei Yang, Hao Lin, Zhenhua Li, and Feng Qian. Fusing speed index during Web page loading. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 23:1–23:23, March 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3511214.

Lin:2019:COO

[LYP^{+}19] Qiulin Lin, Hanling Yi, John Pang, Minghua Chen, Adam Wierman, Michael Honig, and Yuanzhang Xiao. Competitive online optimization under inventory constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):10:1–10:28, March 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3322205.
3311081.

Li:2022:RCL

[LYQ^{+}22] Tongxin Li, Ruixiao Yang, Guannan Qu, Guanya Shi, Chenkai Yu, Adam Wierman, and Steven Low. Robustness and consistency in linear quadratic control with untrusted predictions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 18:1–18:35, March 2022. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3508038.

Li:2022:ECL

[LYR22] Pengfei Li, Jianyi Yang, and Shaolei Ren. Expert-calibrated learning for online optimization with switching costs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2): 28:1–28:??, June 2022. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3530894.

Lu:2023:DMA

[LZC^{+}23] Haoran Lu, Qingchuan Zhao, Yongliang Chen, Xiaojing Liao, and Zhiqiang Lin. Detecting and measuring aggressive location harvesting in mobile apps via data-flow path embedding. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):18:1– 18:??, March 2023. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3579447.

Liu:2022:LPM

[LZGN22] Yilin Liu, Shijia Zhang, Mahanth Gowda, and Srihari Nelakuditi. Leveraging the properties of mmWave signals for 3D finger motion tracking for interactive IoT applications. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(3):52:1–52:??, December 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3570613.

Li:2023:PKB

[LZJ23] Fengjiao Li, Xingyu Zhou, and Bo Ji. (Private) kernelized bandits with distributed biased feedback. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1): 5:1–5:??, March 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3579318.

Mizrahi:2023:IBL

[MBLYR23] Avi Mizrahi, Daniella Bar-Lev, Eitan Yaakobi, and Ori Rottenstreich. Invertible Bloom lookup tables with listing guarantees. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(3):61:1– 61:??, December 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3626792.

Mukherjee:2018:AOL

[MBvL18] Debankur Mukherjee, Sem C.

Borst, and Johan S. H. van Leeuwaarden. Asymptotically optimal load balancing topologies. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):14:1–14:29, April 2018. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3179417.

Mallick:2019:RCN

[MCS^{+}19] Ankur Mallick, Malhar Chaudhari, Utsav Sheth, Ganesh Palanikumar, and Gauri Joshi. Rateless codes for near-perfect load balancing in distributed matrix–vector multiplication. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 3(3):58:1–58:40, December 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3366706.

Mukherjee:2017:OSE

[MDBvL17] Debankur Mukherjee, Souvik Dhara, Sem C. Borst, and Johan S. H. van Leeuwaarden. Optimal service elasticity in large-scale distributed systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):25:1– 25:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084463.

Moura:2024:DDN

[MDS^{+}24] Giovane C. M. Moura, Marco Davids, Caspar Schutijser, Cristian Hesselman, John Heidemann, and Georgios Smaragdakis. Deep dive into NTP Pool’s popularity and mapping. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(1):15:1–15:??, March 2024. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3639041.

Marder:2020:VFO

[MLHC20] Alexander Marder, Matthew Luckie, Bradley Huffaker, and KC Claffy. vrfinder: Finding outbound addresses in Traceroute. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):40:1–40:28, June 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3392158.

Monterubbiano:2023:LAR

[MLW^{+}23] Andrea Monterubbiano, Jonatan Langlet, Stefan Walzer, Gianni Antichi, Pedro Reviriego, and Salvatore Pontarelli. Lightweight acquisition and ranging of flows in the data plane. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7 (3):44:1–44:??, December 2023. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3626775.

MacMillan:2023:CAO

[MMS^{+}23] Kyle MacMillan, Tarun Mangla, James Saxon, Nicole P. Marwell, and Nick Feamster. A comparative analysis of ookla speedtest and measurement labs network diagnostic test (NDT7). Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):19:1–19:??, March 2023. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3579448.

Magureanu:2017:OLO

[MPIZ17] Stefan Magureanu, Alexandre Proutiere, Marcus Isaksson, and Boxun Zhang. Online learning of optimally diverse rankings. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):32:1– 32:26, December 2017. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3154490.

Mishra:2019:GIT

[MSJ^{+}19] Ayush Mishra, Xiangpeng Sun, Atishya Jain, Sameer Pande, Raj Joshi, and Ben Leong. The Great Internet TCP Congestion Control Census. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):45:1–45:24, December 2019. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3366693.

Mhaisen:2022:ONR

[MSPI22] Naram Mhaisen, Abhishek Sinha, Georgios Paschos, and George Iosifidis. Optimistic no-regret algorithms for discrete caching. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(3):48:1–48:??, December 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3570608.

Mahgoub:2022:WWC

[MYS^{+}22] Ashraf Mahgoub, Edgardo Barsallo Yi, Karthick Shankar, Eshaan Minocha, Sameh Elnikety, Saurabh Bagchi, and Somali Chaterji. WISEFUSE: Workload characterization and DAG transformation for serverless workflows. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2):26:1– 26:??, June 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3530892.

Naseer:2023:JCF

[NB23] Usama Naseer and Theophilus A.

Benson. JS capsules: a framework for capturing fine-grained JavaScript memory measurements for the mobile Web. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):14:1–14:??, March 2023. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3579327.

Ngoc:2019:EYS

[NBB^{+}19] Tu Dinh Ngoc, Bao Bui, Stella Bitchebe, Alain Tchana, Valerio Schiavoni, Pascal Felber, and Daniel Hagimont. Everything you should know about Intel SGX performance on virtualized systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1): 5:1–5:21, March 2019. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3322205.
3311076.

Nam:2021:XRN

[NGB^{+}21] Yun Seong Nam, Jianfei Gao, Chandan Bothra, Ehab Ghabashneh, Sanjay Rao, Bruno Ribeiro, Jibin Zhan, and Hui Zhang. Xatu: Richer neural network based prediction for video streaming. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):44:1–44:26, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491056.

Nitu:2018:WSS

[NKY^{+}18] Vlad Nitu, Aram Kocharyan, Hannas Yaya, Alain Tchana, Daniel Hagimont, and Hrachya Astsatryan. Working set size estimation techniques in virtualized environments: One size does not fit all. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1): 19:1–19:22, April 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179422.

Nguyen:2017:OIC

[NNVD17] Hung T. Nguyen, Tri P.

Nguyen, Tam N. Vu, and Thang N. Dinh. Outward influence and cascade size estimation in billion-scale networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):20:1– 20:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084457.

Nikolopoulos:2019:RPS

[NPAP19] Pavlos Nikolopoulos, Christos Pappas, Katerina Argyraki, and Adrian Perrig. Retroactive packet sampling for traffic receipts. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):19:1–19:39, March 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3322205.3311090.

Nguyen:2019:NRA

[NT19] Lan N. Nguyen and My T.

Thai. Network resilience assessment via QoS degradation metrics: an algorithmic approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 3(1):1:1–1:32, March 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3322205.3311072.

Nain:2020:AME

[NVGT20] Philippe Nain, Gayane Vardoyan, Saikat Guha, and Don Towsley. On the analysis of a multipartite entanglement distribution switch. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2): 23:1–23:39, June 2020. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3392141.

Ortiz:2024:SWE

[OCW24] Neil Ortiz, Alvaro A. Cardenas, and Avishai Wool. SCADA world: an exploration of the diversity in power grid networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(1):10:1– 10:??, March 2024. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3639036.

Oleksenko:2018:IME

[OKB^{+}18] Oleksii Oleksenko, Dmitrii Kuvaiskii, Pramod Bhatotia, Pascal Felber, and Christof Fetzer. Intel MPX explained: a cross-layer analysis of the Intel MPX system stack. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (2):28:1–28:30, June 2018. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3224423.

Pourghassemi:2021:ACP

[PBLC21] Behnam Pourghassemi, Jordan Bonecutter, Zhou Li, and Aparna Chandramowlishwaran. adPerf: Characterizing the performance of thirdparty ads. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(1):03:1–03:26, February 2021. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3447381.

Psychas:2017:NPV

[PG17] Konstantinos Psychas and Javad Ghaderi. On nonpreemptive VM scheduling in the cloud. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):35:1–35:29, December 2017. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3154493.

Perivier:2021:RTA

[PHSB21] Noémie Périvier, Chamsi Hssaine, Samitha Samaranayake, and Siddhartha Banerjee. Realtime approximate routing for smart transit systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2): 24:1–24:30, June 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3460091.

Pi:2020:LIA

[PJDQ20] Yibo Pi, Sugih Jamin, Peter Danzig, and Feng Qian. Latency imbalance among Internet load-balanced paths: a cloud-centric view. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2): 32:1–32:29, June 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392150.

Pan:2022:FLC

[PLX22] Yueyang Pan, Ruihan Li, and Chenren Xu. The first 5G-LTE comparative study in extreme mobility. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1):20:1– 20:22, March 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508040.

Pachilakis:2021:YMA

[PPL^{+}21] Michalis Pachilakis, Panagiotis Papadopoulos, Nikolaos Laoutaris, Evangelos P. Markatos, and Nicolas Kourtellis. YourAdvalue: Measuring advertising price dynamics without bankrupting user privacy. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 5(3):32:1–32:26, December 2021. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3491044.

Pourghassemi:2019:WIA

[PSC19] Behnam Pourghassemi, Ardalan Amiri Sani, and Aparna Chandramowlishwaran. What-if analysis of page load time in Web browsers using causal profiling. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):27:1– 27:23, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326142.

Pan:2022:OOF

[PSLW22] Weici Pan, Guanya Shi, Yiheng Lin, and Adam Wierman. Online optimization with feedback delay and nonlinear switching cost. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 17:1–17:34, March 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508037.

Pan:2023:SRE

[PSS23] Jiayu Pan, Yin Sun, and Ness B. Shroff. Sampling for remote estimation of the Wiener process over an unreliable channel. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7 (3):60:1–60:??, December 2023. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3626791.

Pignolet:2017:TNP

[PST17] Yvonne-Anne Pignolet, Stefan Schmid, and Gilles Tredan. Tomographic node placement strategies and the impact of the routing model. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):42:1–42:23, December 2017. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3154501.

Pyrgelis:2020:MMP

[PTD20] Apostolos Pyrgelis, Carmela Troncoso, and Emiliano De Cristofaro. Measuring membership privacy on aggregate location time-series. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2): 36:1–36:28, June 2020. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3392154.

Pandey:2024:TRT

[PYL24] Santosh Pandey, Amir Yazdanbakhsh, and Hang Liu. TAO: Re-thinking DL-based microarchitecture simulation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(2): 28:1–28:??, June 2024. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3656012.

Pan:2023:HPT

[PZB^{+}23] Yicheng Pan, Yang Zhang, Tingzhu Bi, Linlin Han, Yu Zhang, Meng Ma, Xiangzhuang Shen, Xinrui Jiang, Feng Wang, Xian Liu, and Ping Wang. HEAL: Performance troubleshooting deep inside data center hosts. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7 (3):54:1–54:??, December 2023. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3626785.

Quan:2019:NFM

[QTES19] Guocong Quan, Jian Tan, Atilla Eryilmaz, and Ness Shroff. A new flexible multiflow LRU cache management paradigm for minimizing misses. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):39:1– 39:30, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326154.

Quach:2017:ILT

[QWQ17] Alan Quach, Zhongjie Wang, and Zhiyun Qian. Investigation of the 2016 Linux TCP stack vulnerability at scale. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):4:1– 4:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

                                        /dl.acm.org/citation.cfm?
                                        id=3084441.

Raaijmakers:2020:ASR

[RB20] Youri Raaijmakers and Sem Borst. Achievable stability in redundancy systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):46:1–46:21, November 2020. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3428331.

Randone:2021:RMF

[RBT21] Francesca Randone, Luca Bortolussi, and Mirco Tribastone. Refining mean-field approximations by dynamic state truncation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2):25:1– 25:30, June 2021. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3460092.

Rutten:2023:SOO

[RCMW23] Daan Rutten, Nicolas Christianson, Debankur Mukherjee, and Adam Wierman. Smoothed online optimization with unreliable predictions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):12:1–12:??, March 2023. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3579442.

Ros-Giralt:2019:BSC

[RGBY^{+}19] Jordi Ros-Giralt, Atul Bohara, Sruthi Yellamraju, M. Harper Langston, Richard Lethin, Yuang Jiang, Leandros Tassiulas, Josie Li, Yuanlong Tan, and Malathi Veeraraghavan. On the bottleneck structure of congestion-controlled networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 3(3):59:1–59:31, December 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3366707.

Ramanathan:2023:NFC

[RJ23] Alagappan Ramanathan and Sangeetha Abdu Jyothi. Nautilus: a framework for crosslayer cartography of submarine cables and IP links. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7 (3):46:1–46:??, December 2023. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3626777.

Ryoo:2022:MSR

[RKK22] Jihyun Ryoo, Mahmut Taylan Kandemir, and Mustafa Karakoy. Memory space recycling. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1):14:1– 14:24, March 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508034.

Rana:2022:FAR

[RKTV22] Ranvir Rana, Sreeram Kannan, David Tse, and Pramod Viswanath. Free2Shard: Adversary-resistant distributed resource allocation for blockchains. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):11:1–11:38, March 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3508031.

Regmi:2022:APM

[RS22] Hem Regmi and Sanjib Sur. Argus: Predictable millimeterwave picocells with vision and learning augmentation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (1):2:1–2:26, March 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508022.

Ramanathan:2024:XIC

[RSJ24] Alagappan Ramanathan, Rishika Sankaran, and Sangeetha Abdu Jyothi. Xaminer: an Internet cross-layer resilience analysis tool. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(1):16:1–16:??, March 2024. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3639042.

Roy:2024:SMB

[RT24] Rohan Basu Roy and Devesh Tiwari. StarShip: Mitigating I/O bottlenecks in serverless computing for scientific workflows. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 8(1):2:1–2:??, March 2024. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3639028.

Radulovic:2019:PMS

[RVC^{+}19] Milan Radulovic, Rommel Sánchez Verdejo, Paul Carpenter, Petar Radojković, Bruce Jacob, and Eduard Ayguadé. PROFET: Modeling system performance and energy without simulating the CPU. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):34:1– 34:33, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326149.

Salamatian:2022:CBA

[SAM^{+}22] Loqman Salamatian, Scott Anderson, Joshua Matthews, Paul Barford, Walter Willinger, and Mark Crovella. Curvaturebased analysis of network connectivity in private backbone infrastructures. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 5:1–5:32, March 2022. CO-DEN ???? ISSN 2476-1249.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3508025.

Suh:2022:CES

[SATN22] Young-Kyoon Suh, Junyoung An, Byungchul Tak, and Gap-Joo Na. A comprehensive empirical study of query performance across GPU DBM-Ses. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):4:1–4:29, March 2022. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3508024.

Sinclair:2019:ADE

[SBY19] Sean R. Sinclair, Siddhartha Banerjee, and Christina Lee Yu. Adaptive discretization for episodic reinforcement learning in metric spaces. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):55:1–55:44, December 2019. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3366703.

Shi:2024:CLI

[SCH^{+}24] Ruizhe Shi, Ruizhi Cheng, Bo Han, Yue Cheng, and Songqing Chen. A closer look into IPFS: Accessibility, content, and performance. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8 (2):20:1–20:??, June 2024. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3656015.

Senen-Cerda:2022:ACR

[SCS22] Albert Senen-Cerda and Jaron Sanders. Asymptotic convergence rate of dropout on shallow linear neural networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2):32:1– 32:??, June 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3530898.

Sam:2023:OLH

[SCY23] Tyler Sam, Yudong Chen, and Christina Lee Yu. Overcoming the long horizon barrier for sample-efficient reinforcement learning with latent low-rank structure. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(2): 29:1–29:??, June 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3589973.

Shin:2021:PBP

[SCYO21] Suho Shin, Hoyong Choi, Yung Yi, and Jungseul Ok. Power of bonus in pricing for crowdsourcing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):36:1– 36:25, December 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491048.

Subramanian:2018:SFT

[SDA18] Kausik Subramanian, Loris D’Antoni, and Aditya Akella. Synthesis of fault-tolerant distributed router configurations. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(1):22:1–22:26, April 2018. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3179425.

Schardl:2017:CFC

[SDD^{+}17] Tao B. Schardl, Tyler Denniston, Damon Doucet, Bradley C. Kuszmaul, I-Ting Angelina Lee, and Charles E. Leiserson. The CSI framework for compiler-inserted program instrumentation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):43:1–43:25, December 2017. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3154502.

Scully:2020:GPN

[SGHB20] Ziv Scully, Isaac Grosof, and Mor Harchol-Balter. The Gittins policy is nearly optimal in the M/G/k under extremely general conditions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):43:1–43:29, November 2020. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3428328.

Sankararaman:2019:SLM

[SGS19] Abishek Sankararaman, Ayalvadi Ganesh, and Sanjay Shakkottai. Social learning in multi agent multi armed bandits. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):53:1– 53:35, December 2019. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366701.

Scully:2018:SOC

[SHBSW18] Ziv Scully, Mor Harchol-Balter, and Alan Scheller-Wolf. SOAP: One clean analysis of all agebased scheduling policies. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (1):16:1–16:30, April 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179419.

Scully:2020:SNO

[SHBSW20] Ziv Scully, Mor Harchol-Balter, and Alan Scheller-Wolf. Simple near-optimal scheduling for the M/G/1. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):11:1– 11:29, May 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3379477.

Salem:2022:ELT

[SIN22] Tareq Si Salem, Georgios Iosifidis, and Giovanni Neglia. Enabling long-term fairness in dynamic resource allocation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (3):46:1–46:??, December 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3570606.

Shao:2020:YNM

[SIR20] Zhihui Shao, Mohammad A. Islam, and Shaolei Ren. Your noise, my signal: Exploiting switching noise for stealthy data exfiltration from desktop computers. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):07:1– 07:39, May 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3379473.

Sharma:2017:PDR

[SIS17] Prateek Sharma, David Irwin, and Prashant Shenoy. Portfolio-driven resource management for transient cloud servers. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):5:1– 5:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084442.

Sermpezis:2019:ICI

[SK19] Pavlos Sermpezis and Vasileios Kotronis. Inferring catchment in Internet routing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2): 30:1–30:31, June 2019. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3341617.
3326145.

Shi:2019:VLA

[SLJ19] Ming Shi, Xiaojun Lin, and Lei Jiao. On the value of look-ahead in competitive online convex optimization. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (2):22:1–22:42, June 2019. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3341617.
3326136.

Song:2020:UCS

[SLWS20] JinKe Song, Qiang Li, Haining Wang, and Limin Sun. Under the concealing surface: Detecting and understanding live webcams in the wild. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1): 05:1–05:25, May 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3379471.

Silva:2022:AIB

[SMF^{+}22] Brivaldo A. Silva, Paulo Mol, Osvaldo Fonseca, Italo Cunha, Ronaldo A. Ferreira, and Ethan Katz-Bassett. Automatic inference of BGP location communities. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):3:1–3:23, March 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3508023.

Shi:2020:LHC

[SQ20] Shouqian Shi and Chen Qian. Ludo Hashing: Compact, fast, and dynamic key-value lookups for practical network systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):22:1– 22:32, June 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3392140.

Sejourne:2018:PFM

[SSB18] Thibault Séjournè, Samitha Samaranayake, and Siddhartha Banerjee. The price of fragmentation in mobility-on-demand services. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2):30:1– 30:26, June 2018. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3224425.

Sun:2017:ASM

[SSM^{+}17] Wen Sun, Veronique Simon, Sebastien Monnet, Philippe Robert, and Pierre Sens. Analysis of a stochastic model of replication in large distributed storage systems: A mean-field approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):24:1– 24:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084462.

Singh:2021:PNP

[STGM21] Rachee Singh, David Tench, Phillipa Gill, and Andrew Mc-Gregor. PredictRoute: a network path prediction toolkit. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 5(2):23:1–23:24, June 2021. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3460090.

Scully:2020:CPO

[SvKB^{+}20] Ziv Scully, Lucas van Kreveld, Onno Boxma, Jan-Pieter Dorsman, and Adam Wierman. Characterizing policies with optimal response time tails under heavy-tailed job sizes. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (2):30:1–30:33, June 2020. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3392148.

Snyder:2020:WFF

[SVL20] Peter Snyder, Antoine Vastel, and Ben Livshits. Who filters the filters: Understanding the growth, usefulness and efficiency of crowdsourced ad blocking. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):26:1– 26:24, June 2020. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3392144.

Su:2019:SDG

[SX19] Lili Su and Jiaming Xu. Securing distributed gradient descent in high dimensional statistical learning. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):12:1–12:41, March 2019. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3322205.3311083.

Sun:2022:OKP

[SYH^{+}22] Bo Sun, Lin Yang, Mohammad Hajiesmaili, Adam Wierman, John C. S. Lui, Don Towsley, and Danny H. K. Tsang. The online knapsack problem with departures. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (3):57:1–57:??, December 2022. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3570618.

Su:2019:CLB

[SZL19] Lili Su, Martin Zubeldia, and Nancy Lynch. Collaboratively learning the best option on graphs, using bounded local memory. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):11:1–11:32, March 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3322205.3311082.

Sun:2020:CAO

[SZL^{+}20] Bo Sun, Ali Zeynali, Tongxin Li, Mohammad Hajiesmaili, Adam Wierman, and Danny H. K. Tsang. Competitive algorithms for the online multiple knapsack problem with application to electric vehicle charging. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(3):51:1– 51:32, November 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3428336.

Tsanikidis:2023:NOP

[TG23] Christos Tsanikidis and Javad Ghaderi. Near-optimal packet scheduling in multihop networks with end-to-end deadline constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(3):50:1– 50:??, December 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3626781.

Tang:2023:SLR

[TKFJ23] Weizhao Tang, Lucianna Kiffer, Giulia Fanti, and Ari Juels. Strategic latency reduction in blockchain peerto-peer networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(2): 32:1–32:??, June 2023. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3589976.

Tang:2021:MMR

[TKK21] Xulong Tang, Mahmut Taylan Kandemir, and Mustafa Karakoy. Mix and match: Reorganizing tasks for enhancing data locality. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2): 20:1–20:24, June 2021. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3460087.

Tang:2018:CND

[TKZ^{+}18] Xulong Tang, Mahmut Taylan Kandemir, Hui Zhao, Myoungsoo Jung, and Mustafa Karakoy. Computing with near data. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(3):42:1– 42:30, December 2018. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3287321.

Thomas:2020:TSI

[TL20] Ludovic Thomas and Jean-Yves Le Boudec. On time synchronization issues in timesensitive networks with regulators and nonideal clocks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (2):27:1–27:41, June 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392145.

Tan:2018:SMV

[TLL^{+}18] Zhaowei Tan, Yuanjie Li, Qianru Li, Zhehui Zhang, Zhehan Li, and Songwu Lu. Supporting mobile VR in LTE networks: How close are we? Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (1):8:1–8:31, April 2018. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3179411.

Tay:2020:E

[TM20] Y. C. Tay and Athina Markopoulou.

Editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):21:1, June 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3392139.

Tillberg:2020:OBS

[TMM20] Erik Tillberg, Peter Marbach, and Ravi Mazumdar. Optimal bidding strategies for online ad auctions with overlapping targeting criteria. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (2):34:1–34:55, June 2020. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3392152.

Talebi:2018:LPF

[TP18] Mohammad Sadegh Talebi and Alexandre Proutiere. Learning proportionally fair allocations with low regret. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2): 36:1–36:31, June 2018. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3224431.

Tang:2018:QDL

[TPK^{+}18] Xulong Tang, Ashutosh Pattnaik, Onur Kayiran, Adwait Jog, Mahmut Taylan Kandemir, and Chita Das. Quantifying data locality in dynamic parallelism in GPUs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (3):39:1–39:24, December 2018. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3287318.

Tan:2018:RPS

[TQJS18] Jian Tan, Guocong Quan, Kaiyi Ji, and Ness Shroff. On resource pooling and separation for LRU caching. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (1):5:1–5:31, April 2018. CO-DEN ???? ISSN 2476-1249.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3179408.

Tang:2019:RWB

[TS19] Dengwang Tang and Vijay G.

Subramanian. Random walk based sampling for load balancing in multi-server systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):14:1–14:44, March 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3322205.3311085.

Tan:2020:MDO

[TSLG^{+}20] Xiaoqi Tan, Bo Sun, Alberto Leon-Garcia, Yuan Wu, and Danny H. K. Tsang. Mechanism design for online resource allocation: a unified approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):24:1– 24:46, June 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3392142.

Tang:2021:RMG

[TTL^{+}21] Jing Tang, Xueyan Tang, Andrew Lim, Kai Han, Chongshou Li, and Junsong Yuan. Revisiting modified greedy algorithm for monotone submodular maximization with a knapsack constraint. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (1):08:1–08:22, February 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3447386.

Tang:2020:PUT

[TWFO20] Weizhao Tang, Weina Wang, Giulia Fanti, and Sewoong Oh. Privacy-utility tradeoffs in routing cryptocurrency over payment channel networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):29:1– 29:39, June 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3392147.

VanHoudt:2019:GAO

[Van19] Benny Van Houdt. Global attraction of ODE-based mean field models with hyperexponential job sizes. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):23:1– 23:23, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326137.

VanHoudt:2022:SAI

[Van22] Benny Van Houdt. On the stochastic and asymptotic improvement of first-come firstserved and nudge scheduling. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(3):50:1–50:??, December 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3570610.

vanderBoor:2019:HSJ

[vdBBvL19] Mark van der Boor, Sem Borst, and Johan van Leeuwaarden. Hyper-scalable JSQ with sparse feedback. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1): 4:1–4:37, March 2019. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3322205.
3311075.

Venkatakrishnan:2017:DRB

[VFV17] Shaileshh Bojja Venkatakrishnan, Giulia Fanti, and Pramod Viswanath. Dandelion: Redesigning the Bitcoin network for anonymity. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):22:1– 22:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084459.

Vlachou:2018:LTL

[VPK18] Christina Vlachou, Ioannis Pefkianakis, and Kyu-Han Kim. LTERadar: Towards LTEaware Wi-Fi access points. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (2):33:1–33:31, June 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3224428.

Vial:2019:SRP

[VS19] Daniel Vial and Vijay Subramanian. A structural result for personalized PageRank and its algorithmic consequences. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):25:1– 25:88, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326140.

Vial:2022:RMA

[VSS22] Daniel Vial, Sanjay Shakkottai, and R. Srikant. Robust multi-agent bandits over undirected graphs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (3):53:1–53:??, December 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3570614.

Wierman:2018:ME

[WA18] Adam Wierman and Aditya Akella. Message from the Editors. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(2):23:1, June 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3224418.

Wang:2018:NNM

[WCX^{+}18] Mowei Wang, Yong Cui, Shihan Xiao, Xin Wang, Dan Yang, Kai Chen, and Jun Zhu. Neural network meets DCN: Traffic-driven topology adaptation with deep learning. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (2):26:1–26:25, June 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3224421.

Williams:2022:MMD

[WHBW22] Jalani K. Williams, Mor Harchol-Balter, and Weina Wang. The M/M/k with deterministic setup times. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (3):56:1–56:??, December 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3570617.

Wei:2019:QMO

[WKWY19] Honghao Wei, Xiaohan Kang, Weina Wang, and Lei Ying. QuickStop: a Markov optimal stopping approach for quickest misinformation detection. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):41:1– 41:25, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326156.

Wang:2017:CMP

[WLdA^{+}17] Brandon Wang, Xiaoye Li, Leandro P. de Aguiar, Daniel S. Menasche, and Zubair Shafiq. Characterizing and modeling patching practices of industrial control systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):18:1– 18:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

                                        /dl.acm.org/citation.cfm?
                                        id=3084455.

Wang:2019:FLA

[WLS19] Sinong Wang, Jiashang Liu, and Ness Shroff. Fundamental limits of approximate gradient coding. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):52:1– 52:22, December 2019. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3366700.

Wang:2022:CCT

[WLZ^{+}22] Kailong Wang, Yuxi Ling, Yanjun Zhang, Zhou Yu, Haoyu Wang, Guangdong Bai, Beng Chin Ooi, and Jin Song Dong. Characterizing cryptocurrency-themed malicious browser extensions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (3):43:1–43:??, December 2022. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3570603.

Wang:2020:PPA

[WM20] Xin Wang and Richard T. B. Ma. On private peering agreements between content and access providers: a contractual equilibrium analysis. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):41:1–41:32, November 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3428326.

Wang:2023:RTS

[WMMC23] Haibo Wang, Dimitrios Melissourgos, Chaoyi Ma, and Shigang Chen. Real-time spread burst detection in data streaming. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(2):35:1–35:??, June 2023. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3589979.

Wang:2017:OTS

[WMX17] Xin Wang, Richard T. B. Ma, and Yinlong Xu. On optimal two-sided pricing of congested networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):7:1– 7:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084444.

Wang:2017:SGN

[WS17a] Sinong Wang and Ness Shroff.

Security game with nonadditive utilities and multiple attacker resources. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):

13:1–13:??, June 2017. CO-DEN ???? ISSN 2476-1249.

URL  http://dl.acm.org/
citation.cfm?id=3084450.

Wang:2017:TFC

[WS17b] Sinong Wang and Ness Shroff. Towards fast-convergence, lowdelay and low-complexity network optimization. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):34:1–34:32, December 2017. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154492.

Wang:2024:BNO

[WSC^{+}24] Lingdong Wang, Simran Singh, Jacob Chakareski, Mohammad Hajiesmaili, and Ramesh K. Sitaraman. BONES: Nearoptimal neural-enhanced video streaming. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(2):19:1– 19:??, June 2024. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3656014.

Wang:2017:UBI

[WUNK17] Cheng Wang, Bhuvan Urgaonkar, Neda Nasiriani, and George Kesidis. Using burstable instances in the public cloud: Why, when and how? Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (1):11:1–11:??, June 2017. CO-DEN ???? ISSN 2476-1249.

URL  http://dl.acm.org/
citation.cfm?id=3084448.

Weng:2020:AZA

[WW20] Wentao Weng and Weina Wang. Achieving zero asymptotic queueing delay for parallel jobs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(3):42:1–42:36, November 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3428327.

Wang:2022:MDA

[WWH^{+}22] Liu Wang, Haoyu Wang, Ren He, Ran Tao, Guozhu Meng, Xiapu Luo, and Xuanzhe Liu. MalRadar: Demystifying Android malware in the new era. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2):40:1–40:??, June 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3530906.

Wang:2021:ZQM

[WXHB21] Weina Wang, Qiaomin Xie, and Mor Harchol-Balter. Zero queueing for multi-server jobs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 5(1):07:1–07:25, February 2021. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3447385.

Wang:2022:UDC

[WXW22] Minhu Wang, Mingwei Xu, and Jianping Wu. Understanding I/O direct cache access performance for end host networking. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1):22:1– 22:37, March 2022. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3508042.

Wei:2018:OLW

[WYN18] Xiaohan Wei, Hao Yu, and Michael J. Neely. Online learning in weakly coupled Markov decision processes: a convergence time study. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1): 12:1–12:38, April 2018. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3179415.

Wei:2020:OPD

[WYN20] Xiaohan Wei, Hao Yu, and Michael J. Neely. Online primal-dual mirror descent under stochastic constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (2):39:1–39:36, June 2020. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3392157.

Weng:2020:OLB

[WZS20] Wentao Weng, Xingyu Zhou, and R. Srikant. Optimal load balancing with locality constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(3):45:1– 45:37, November 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3428330.

Wei:2019:HBS

[WZT19] Song Wei, Kun Zhang, and Bibo Tu. HyperBench: a benchmark suite for virtualization capabilities. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):24:1– 24:22, June 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3341617.3326138.

Xie:2024:HTO

[XGS24] Runhan Xie, Isaac Grosof, and Ziv Scully. Heavytraffic optimal size- and stateaware dispatching. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(1): 9:1–9:??, March 2024. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3639035.

Xie:2022:NSF

[XJ22] Yaxiong Xie and Kyle Jamieson.

NG-Scope: Fine-grained telemetry for NextG cellular networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1):12:1–12:26, March 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3508032.

Xiao:2022:MTD

[XLY^{+}22] Dongwei Xiao, Zhibo LIU, Yuanyuan Yuan, Qi Pang, and Shuai Wang. Metamorphic testing of deep learning compilers. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1):15:1– 15:28, March 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508035.

Xiao:2022:MSB

[XVK22] Yunming Xiao, Matteo Varvello, and Aleksandar Kuzmanovic. Monetizing spare bandwidth: The case of distributed VPNs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(2):33:1–33:??, June 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3530899.

Xia:2021:TTD

[XWG^{+}21] Pengcheng Xia, Haoyu Wang, Bingyu Gao, Weihang Su, Zhou Yu, Xiapu Luo, Chao Zhang, Xusheng Xiao, and Guoai Xu. Trade or trick?: Detecting and characterizing scam tokens on Uniswap decentralized exchange. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):39:1– 39:26, December 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491051.

Yang:2022:HLA

[YCH^{+}22] Lin Yang, Yu-Zhen Janice Chen, Mohammad H. Hajiesmaili, Mark Herbster, and Don Towsley. Hierarchical learning algorithms for multi-scale expert problems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2):34:1– 34:??, June 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3530900.

Yang:2019:AER

[YCX^{+}19] Zhenjie Yang, Yong Cui, Shihan Xiao, Xin Wang, Minming Li, Chuming Li, and Yadong Liu. Achieving efficient routing in reconfigurable DCNs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):47:1–47:30, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366695.

Yang:2018:OAO

[YDH^{+}18] Lin Yang, Lei Deng, Mohammad H. Hajiesmaili, Cheng Tan, and Wing Shing Wong. An optimal algorithm for online non-convex learning. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (2):25:1–25:25, June 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3224420.

Yang:2017:PIA

[YHG^{+}17] Sen Yang, Yan He, Zihui Ge, Dongmei Wang, and Jun Xu. Predictive impact analysis for designing a resilient cellular backhaul network. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):30:1–30:33, December 2017. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154488.

Yang:2020:OLO

[YHS^{+}20] Lin Yang, Mohammad H. Hajiesmaili, Ramesh Sitaraman, Adam Wierman, Enrique Mallada, and Wing S. Wong. Online linear optimization with inventory management constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):16:1– 16:29, May 2020. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3379482.

Yu:2023:OAS

[YHW23] Jing Yu, Dimitar Ho, and Adam Wierman. Online adversarial stabilization of unknown networked systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1): 26:1–26:??, March 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3579452.

Ying:2017:SMM

[Yin17] Lei Ying. Stein’s method for mean field approximations in light and heavy traffic regimes. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):12:1– 12:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084449.

Yang:2024:OAR

[YLIR24] Jianyi Yang, Pengfei Li, Mohammad Jaminur Islam, and Shaolei Ren. Online allocation with replenishable budgets: Worst case and beyond. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8 (1):4:1–4:??, March 2024. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3639030.

Yang:2017:SRL

[YLX17] Sen Yang, Bill Lin, and Jun Xu. Safe randomized loadbalanced switching by diffusing extra loads. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):29:1–29:37, December 2017. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3154487.

Yang:2021:SSG

[YNJS21] Lishan Yang, Bin Nie, Adwait Jog, and Evgenia Smirni. SUGAR: Speeding up GPGPU application resilience estimation with input sizing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (1):01:1–01:29, February 2021. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3447375.

Yun:2018:MAB

[YPA^{+}18] Donggyu Yun, Alexandre Proutiere, Sumyeong Ahn, Jinwoo Shin, and Yung Yi. Multiarmed bandit with additional observations. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):13:1–13:22, April 2018. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3179416.

Yu:2024:STO

[YS24] George Yu and Ziv Scully. Strongly tail-optimal scheduling in the light-tailed M/G/1. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 8(2):27:1–27:??, June 2024. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3656011.

Yang:2022:PRC

[YSL^{+}22] Gyeongsik Yang, Changyong Shin, Jeunghwan Lee, Yeonho Yoo, and Chuck Yoo. Prediction of the resource consumption of distributed deep learning systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2): 29:1–29:??, June 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3530895.

Yu:2019:ALB

[YWB19] Haoran Yu, Ermin Wei, and Randall A. Berry. Analyzing location-based advertising for vehicle service providers using effective resistances. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1): 6:1–6:35, March 2019. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3322205.
3311077.

Yang:2017:ORO

[YWH17] Lin Yang, Wing Shing Wong, and Mohammad H. Hajiesmaili. An optimal randomized online algorithm for QoS buffer management. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):36:1–36:26, December 2017.

CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3154494.

Yu:2022:TCN

[YX22] Christina Lee Yu and Xumei Xi. Tensor completion with nearly linear samples given weak side information. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (2):39:1–39:??, June 2022. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3530905.

Yu:2021:PDH

[YXL21] Liren Yu, Jiaming Xu, and Xiaojun Lin. The power of D-hops in matching powerlaw graphs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2):27:1– 27:43, June 2021. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3460094.

Yang:2021:CAO

[YZH^{+}21] Lin Yang, Ali Zeynali, Mohammad H. Hajiesmaili, Ramesh K. Sitaraman, and Don Towsley. Competitive algorithms for online multidimensional knapsack problems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):30:1– 30:30, December 2021. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3491042.

Zhang:2024:MLS

[ZAA^{+}24] Huaifeng Zhang, Mohannad Alhanahnah, Fahmi Abdulqadir Ahmed, Dyako Fatih, Philipp Leitner, and Ahmed Ali-Eldin. Machine learning systems are bloated and vulnerable. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8 (1):6:1–6:??, March 2024. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3639032.

Zhao:2022:DAC

[ZCM^{+}22] Chenxingyu Zhao, Tapan Chugh, Jaehong Min, Ming Liu, and Arvind Krishnamurthy. Dremel: Adaptive configuration tuning of RocksDB KV-Store. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2):37:1– 37:??, June 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3530903.

Zhao:2021:UPG

[ZCW21] Shizhen Zhao, Peirui Cao, and Xinbing Wang. Understanding the performance guarantee of physical topology design for optical circuit switched data centers. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):42:1– 42:24, December 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491054.

Zerwas:2023:DHT

[ZGB^{+}23] Johannes Zerwas, Csaba Györgyi, Andreas Blenk, Stefan Schmid, and Chen Avin. Duo: a highthroughput reconfigurable datacenter network using local routing and control. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):20:1–20:??, March 2023. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3579449.

Zhu:2019:UNP

[ZGN^{+}19] Xiao Zhu, Yihua Ethan Guo, Ashkan Nikravesh, Feng Qian, and Z. Morley Mao. Understanding the networking performance of Wear OS. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (1):3:1–3:25, March 2019. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3322205.
3311074.

Zhang:2021:CHE

[ZHM^{+}21] Yiguang Zhang, Jessy Xinyi Han, Ilica Mahajan, Priyanjana Bengani, and Augustin Chaintreau. Chasm in hegemony: Explaining and reproducing disparities in homophilous networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2):

16:1–16:38, June 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3460083.

Zhou:2022:DPR

[Zho22] Xingyu Zhou. Differentially private reinforcement learning with linear function approximation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1):8:1– 8:27, March 2022. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3508028.

Zhang:2021:SFI

[ZHZ^{+}21] Qingzhao Zhang, David Ke Hong, Ze Zhang, Qi Alfred Chen, Scott Mahlke, and Z. Morley Mao. A systematic framework to identify violations of scenario-dependent driving rules in autonomous vehicle software. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2):15:1– 15:25, June 2021. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3460082.

Zhang:2020:ODP

[ZKAV20] Lei Zhang, Reza Karimi, Irfan Ahmad, and Ymir Vigfusson. Optimal data placement for heterogeneous cache, memory, and storage systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):

06:1–06:27, May 2020. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3379472.

Zheng:2018:HCL

[ZL18] Pengfei Zheng and Benjamin C. Lee. Hound: Causal learning for datacenter-scale straggler diagnosis. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):17:1–17:36, April 2018. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3179420.

Zhang:2023:DMS

[ZLL^{+}23] Mingming Zhang, Xiang Li, Baojun Liu, Jianyu Lu, Yiming Zhang, Jianjun Chen, Haixin Duan, Shuang Hao, and Xiaofeng Zheng. Detecting and measuring security risks of hosting-based dangling domains. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 7(1):9:1–9:??, March 2023. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3579440.

Zhang:2018:PSF

[ZLM^{+}18] Shenglin Zhang, Ying Liu, Weibin Meng, Zhiling Luo, Jiahao Bu, Sen Yang, Peixian Liang, Dan Pei, Jun Xu, Yuzhi Zhang, Yu Chen, Hui Dong, Xianping Qu, and Lei Song. Pre-Fix: Switch failure prediction in datacenter networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (1):2:1–2:29, April 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179405.

Zhang:2017:OPP

[ZLW17] Zijun Zhang, Zongpeng Li, and Chuan Wu. Optimal posted prices for online cloud resource allocation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):23:1– 23:??, June 2017. CODEN ???? ISSN 2476-1249. URL http:/

/dl.acm.org/citation.cfm?
id=3084460.

Zhu:2023:DBF

[ZLX^{+}23] Wenzhe Zhu, Yongkun Li, Erci Xu, Fei Li, Yinlong Xu, and John C. S. Lui. DiffForward: On balancing forwarding traffic for modern cloud block services via differentiated forwarding. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1): 15:1–15:??, March 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3579444.

Zhang:2024:NFA

[ZMC24] Yiguang Zhang, Reetahan Mukhopadhyay, and Augustin Chaintreau. Network fairness ambivalence: When does social network capital mitigate or amplify unfairness? Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(2):25:1– 25:??, June 2024. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3656017.

Zarchy:2019:ACC

[ZMSS19] Doron Zarchy, Radhika Mittal, Michael Schapira, and Scott Shenker. Axiomatizing congestion control. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):33:1– 33:33, June 2019. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3341617.3326148.

Zhu:2020:CTI

[ZMW^{+}20] Pengxiong Zhu, Keyu Man, Zhongjie Wang, Zhiyun Qian, Roya Ensafi, J. Alex Halderman, and Haixin Duan. Characterizing transnational Internet performance and the Great Bottleneck of China. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1): 13:1–13:23, May 2020. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3379479.

Zhang:2023:GCL

[ZQX^{+}23] Yizhou Zhang, Guannan Qu, Pan Xu, Yiheng Lin, Zaiwei Chen, and Adam Wierman.

Global convergence of localized policy iteration in networked multi-agent reinforcement learning. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7(1):13:1– 13:??, March 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3579443.

Zia:2022:TDW

[ZRC^{+}22] Haris Bin Zia, Aravindh Raman, Ignacio Castro, Ishaku Hassan Anaobi, Emiliano De Cristofaro, Nishanth Sastry, and Gareth Tyson. Toxicity in the decentralized Web and the potential for model sharing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(2):35:1– 35:??, June 2022. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3530901.

Zhang:2023:CCB

[ZSPZ23] Ding Zhang, Panneer Selvam Santhalingam, Parth Pathak, and Zizhan Zheng. CoBF: Coordinated beamforming in dense mmWave networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 7 (2):31:1–31:??, June 2023. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3589975.

Zhou:2018:HTD

[ZTS18] Xingyu Zhou, Jian Tan, and Ness Shroff. Heavy-traffic delay optimality in pull-based load balancing systems: Necessary and sufficient conditions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (3):41:1–41:33, December 2018. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3287323.

Zhang:2021:MSW

[ZTY^{+}21] Yue Zhang, Bayan Turkistani, Allen Yuqing Yang, Chaoshun Zuo, and Zhiqiang Lin. A measurement study of Wechat mini-apps. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2):14:1– 14:25, June 2021. CODEN ???? ISSN 2476-1249. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3460081.

Zubeldia:2020:DOP

[Zub20]  Martin Zubeldia.  Delay-

optimal policies in partial forkjoin systems with redundancy and random slowdowns. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (1):02:1–02:49, May 2020. CO-DEN ???? ISSN 2476-1249.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3379468.

Zhou:2017:DLC

[ZWT^{+}17] Xingyu Zhou, Fei Wu, Jian Tan, Yin Sun, and Ness Shroff. Designing low-complexity heavytraffic delay-optimal load balancing schemes: Theory to algorithms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):39:1–39:30, December 2017. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3154498.

Zhou:2018:DQI

[ZWT^{+}18] Xingyu Zhou, Fei Wu, Jian Tan, Kannan Srinivasan, and Ness Shroff. Degree of queue imbalance: Overcoming the limitation of heavy-traffic delay optimality in load balancing systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):21:1–21:41, April 2018. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3179424.

Zhang:2019:AMD

[ZYH^{+}19] Lei Zhang, Zhemin Yang, Yuyu He, Mingqi Li, Sen Yang, Min Yang, Yuan Zhang, and Zhiyun Qian. App in the middle: Demystify application virtualization in Android and its security threats. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):17:1–17:24, March 2019. CODEN ???? ISSN 2476-1249. URL https:

//dl.acm.org/doi/10.1145/
3322205.3311088.

Zhou:2017:PSM

[ZZCC17] You Zhou, Yian Zhou, Min Chen, and Shigang Chen. Persistent spread measurement for big network data based on register intersection. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1): 15:1–15:??, June 2017. CO-DEN ???? ISSN 2476-1249.

                                        URL  http://dl.acm.org/
                                        citation.cfm?id=3084452.

Zhu:2021:FBG

[ZZLL21] Zhaowei Zhu, Jingxuan Zhu, Ji Liu, and Yang Liu. Federated Bandit: a gossiping approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(1):02:1– 02:29, February 2021. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3447380.

Zhou:2019:GSF

[ZZM^{+}19] You Zhou, Youlin Zhang, Chaoyi Ma, Shigang Chen, and Olufemi O. Odegbile. Generalized sketch families for network traffic measurement. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):51:1–51:34, December 2019. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3366699.